### v1

# Learning Contrastive Feature Representations for Facial Action Unit Detection

Abstract
The predominant approach to facial action unit (AU) detection revolves around a supervised multi-label binary classification problem. Existing methodologies often encode pixel-level information of AUs, thereby imposing substantial demands on model complexity and expressiveness. Moreover, this practice elevates the susceptibility to overfitting due to the presence of noisy AU labels. In the present study, we introduce a contrastive learning framework enhanced by both supervised and self-supervised signals. The objective is to acquire discriminative features, deviating from the conventional pixel-level learning paradigm within the domain of AU detection. To address the challenge posed by noisy AU labels, we augment the supervised signal through the introduction of a self-supervised signal. This augmentation is achieved through positive sample sampling, encompassing three distinct types of positive sample pairs. Furthermore, to mitigate the imbalanced distribution of each AU type, we employ an importance re-weighting strategy tailored for minority AUs. The resulting loss, denoted as AUNCE, is proposed to encapsulate this strategy. Our experimental assessments, conducted on two widely-utilized benchmark datasets (BP4D and DISFA), underscore the superior performance of our approach compared to state-of-the-art methods in the realm of AU detection. Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
Index Terms:
Facial action unit detection, Contrastive learning, Positive sample sampling, Importance re-weighting strategy.
I Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding. It has introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2], enabling their simulation. In contemporary times, AU detection [3, 4, 5, 6] has evolved into a technique utilizing computer vision algorithms to automatically identify and track AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction [7], emotion analysis [8, 9], and monitoring in activities such as car driving [10].
Refer to caption (a)
Refer to caption (b)
Refer to caption (c)
Refer to caption (d)
Figure 1: Illustration of our proposed positive sample sampling strategy. From left to right: (a) original images, (b) self-supervised images obtained by simple augmentations, (c) images with the highest similarity to (a), and (d) images with the lowest similarity to (a). We use (a) with (b), (c), mixture of all positive samples as our positive sample pairs.
In recent years, a predominant trend in supervised frameworks for AU detection has been observed, as evidenced by various studies [11, 12, 13, 14, 15]. These frameworks have predominantly employed the traditional binary cross-entropy loss, which serves as a metric for gauging the disparity between predicted and actual ground truth labels, providing a systematic means to guide the learning process. However, encoding the pixel-level information is notable for its heightened sensitivity to models characterized by a substantial number of parameters, as it endeavors to encode information at the pixel level within the input space. This imposes a notable demand on the fitting capability of the models. Additionally, pixel-level generative learning, while computationally intensive, may prove superfluous for AU detection tasks, given their nature as multi-label binary classification problems. The pursuit of pixel-level features not only introduces unwarranted model complexity but also augments the susceptibility to overfitting.
Representation learning approaches rooted in the discriminative contrastive learning paradigm have demonstrated substantial promise, yielding state-of-the-art outcomes [16, 17, 18, 19, 20]. The fundamental concept of contrastive learning involves maximizing the consistency of positive example features that are semantically similar while minimizing the consistency of negative example features that are semantically dissimilar [21]. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized, relieving the model from encoding pixel-level information [22]. However, the direct integration of contrastive learning into AU detection tasks encounters two principal challenges. The first challenge pertains to the noisy label problem [23]. Due to the weak amplitude of AU occurrence and the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit errors and noisy labels during the labeling process. Consequently, fitting models with noisy and erroneous samples significantly undermines the overall fitting ability of models. The second challenge is the class imbalance problems among different AUs [11]. Most AU classes manifest a substantial imbalance between positive and negative samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance not only diminishes the fitting efficiency of neural networks but also influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class.
To tackle these challenges, we propose a simple contrastive learning framework to encode discriminative features between samples for AU detection, serving as a substitution for conventional pixel-level learning. This approach offers a pathway to lightweight model development within the domain of AU detection tasks. Moreover, in response to the issue of noisy labeling in AUs, we introduce self-supervised signals to augment the noise supervision signal. This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that positive examples derived from data augmentation with the self-supervised signal enhance robustness, as illustrated in Fig. 1. Additionally, to mitigate the class imbalance problem among different types of AUs, we apply re-weighting to the contrastive loss. The resultant loss, resembling the commonly utilized InfoNCE, is termed AUNCE. Consequently, modifications are made to the gradient magnitude during back-propagation in the training process. The primary contributions of our work are succinctly summarized as follows:
  • We propose a novel AUNCE loss to learn contrastive feature representations for AU detection.
  • We introduce a strategy that incorporates self-supervised signal to address the issue of noisy AU labels.
  • We employ an importance re-weighting strategy for minority AUs to tackle the class-imbalance problem.
  • We validate the effectiveness of AUNCE by comparing it to state-of-the-art approaches for AU detection on two widely-used datasets, namely BP4D and DISFA.
II Related Work
We will address two key aspects closely linked to our work: deep contrastive learning and contrastive learning for AU detection.
II-A Deep contrastive learning
Contrastive learning [24] initially emerged from the realm of self-supervised learning, where the self-supervised learning objectives were defined by assigning tasks to agents. Specifically, this approach encompasses the application of diverse image augmentations to identical images, followed by an assessment of the resultant images and their associated features to determine similarity. The primary objective is to augment the similarity among features extracted from the same image while concurrently attenuating the similarity between features derived from distinct images. Subsequently, the evolution of this technology has led to the extension of contrastive learning to supervised and semi-supervised learning paradigms, wherein the utilization of labeled data is leveraged to enhance the representations of model features.
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [25], has witnessed substantial progress. This approach adheres to the learn-to-compare paradigm [26], wherein a distinction is made between observed data and noisy or erroneous data, preventing the model from engaging in the reconstruction of pixel-level information. Despite variations in representation encoders (denoted as $f$) and similarity measures across different tasks [27, 28], the underlying concept remains consistent: training the representation encoder $f$ by optimizing a contrastive loss [29], achieved through contrasting similar pairs ($x,x^{+}$) and dissimilar pairs ($x,x^{-}$).
Prominent frameworks in this domain include SimCLR by Chen et al. [21], which introduces a base encoder network and a projection head trained to maximize agreement utilizing a contrastive loss. Additionally, MoCo, proposed by He et al. [30], employs a momentum network to maintain a queue of negative samples for efficient contrastive learning. Chen and He [31] mitigate undesired collapsing solutions, eschewing the use of negative sample pairs through a stop-gradient operation. Typically, these techniques are founded on the network architecture of quasi-twin neural networks. The image pairs utilized in the training process are derived from the same image, enlarged separately (positive sample pairs). The common practice involves obtaining positive samples through semantic-invariant operations on an anchor, coupled with heavy data augmentation techniques such as random cropping, flipping, image rotation, cutout, and color distortion. Conversely, image pairs are formed by enlarging different images separately (negative sample pairs).
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of numerous positive and negative sample pairs but also incorporates diverse losses for optimization. For example, Oord et al. [22] optimized feature extraction networks by maximizing consistency between predicted and actual outcomes of sequence data, introducing the widely-utilized InfoNCE loss, which implicitly or explicitly serves as a lower bound for mutual information. Theoretical analyses concerning the generalization bounds of contrastive learning for classification tasks have been provided by Arora et al. [32]. Khosla et al. [33] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [34] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
As the technology of contrastive learning advances, there is a discernible shift in the construction of sample pairs for contrastive learning. In cases where data is labeled or partially labeled, positive samples can be constructed using images belonging to the same category, rather than being confined to identical images. This approach augments the diversity of positive sample pairs, consequently enhancing feature expression. Furthermore, contrastive learning serves as a precursor to supervised learning, enabling model pretraining to significantly enhance the performance of subsequent supervised learning models.
II-B Contrastive learning for AU detection
While contrastive learning has garnered widespread popularity in recent years for AU detection, the predominant focus within AU detection methods has been on generative contrastive learning. Various approaches utilize a network with a parallel structure to directly compare features extracted from two or more sampled images. Wu et al. [35] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-affected facial images. Sun et al. [36] proposed an Emotion-aware Contrastive Learning framework designed to capture representations preserving sufficient AU-related information. Zhao et al. [37] devised a contrastive learning scheme aimed at creating a semantic-aware AU manifold, thereby reducing misalignment and lowering the dimensionality of facial action features. Liu et al. [38] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling. This multifaceted approach aimed to enhance the discriminative capabilities of AUs.
Moreover, certain methodologies are specifically designed to incorporate contrastive losses tailored for AU detection tasks, aiming to discern differences among AU classes or different perspectives of the same AU class. Niu et al. [39] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). In a similar vein, Suresh et al. [40] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings. Shang et al. [41] put forth a pixel-level multi-view contrastive loss, enhancing features through distinct cross-scale features. Additionally, Yang et al. [42] suggested a novel weighted multi-label contrastive learning loss tailored for AU detection.
However, AU detection methods grounded in the discriminative contrastive learning paradigm have not found widespread adoption in recent research on AU detection, particularly in the context of unlabeled video sequences. Contemporary endeavors predominantly seek to generalize similarity learning between video frames within the discriminative contrastive learning paradigm, with the objective of acquiring meaningful video-level representations [43, 44, 45, 46, 47, 48]. Nevertheless, these general self-supervised contrastive learning approaches generate representations that may not be specifically tailored to AU detection, leading to debates regarding the existence of a universal representation applicable to all downstream tasks [49]. Despite the inherent temporal coherence and evolution characteristics present in unlabeled video sequences, Li et al. [23] have capitalized on these attributes by formulating a contrastive self-supervised learning method to encode discriminative AU representations from extensive amounts of unlabeled facial videos. However, the practical application of the discriminative contrastive learning paradigm for static images to AU detection frameworks remains theoretical, and the potential substitution of traditional generative cross-entropy loss by such learning methods is a subject of ongoing discussion.
III Method
In the Section. III, we mainly address three main challenges: 1. How to incentivize the encoder to capture differential information among AUs rather than pixel-level information, enabling a lightweight encoder through discriminative contrastive learning paradigm, instead of the traditional cross-entropy classification paradigm. 2. How to solve the class imbalanced problem of each AU type to improve the accuracy of our framework. 3. How to mitigate the problem of AU labels tainted by noise and errors, enhancing the robustness of our models to erroneously annotated data and thereby improving generalization performance of the models.
Figure 2: Overview of the encoder of our contrastive learning framework.
III-A Contrastive Learning
To tackle the first challenge, we propose a contrastive learning framework that encourages the encoder to focus on capturing differences of each AU type instead of the pixel-level information of the whole face. By leveraging the relationships between samples and their corresponding AUs, the framework promotes the extraction of meaningful and discriminative features for AU detection. We investigate the feature representation of a specific AU, denoted as $x$. For AU detection which belongs to multi-label binary classification problem, the cross-entropy loss emerges as the primary choice for training models. This loss captures the discrepancy between predicted and ground truth labels, providing a principled approach to guide the learning process. Specifically, for a specific AU, the classification loss is calculated by comparing the features of samples with their corresponding labels. This enables the model to optimize its parameters with respect to the task of accurately classifying AUs. Mathematically, the typical cross-entropy loss can be expressed as follows:
$\mathcal{L}_{\textsc{ce}}=y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})$ (1)
where $y_{i}$ denotes the ground truth value of occurrence (i.e., 0 or 1) of the $i$-th AU, and $\hat{y}_{i}$ is the corresponding predicted occurrence probability. This loss effectively guides the training process, encouraging the model to allocate higher probabilities to the correct class labels and lower probabilities to incorrect ones. In the case of sample $x_{i}$, the loss aims to minimize the discrepancy between predicted probabilities and the ground truth labels of the models. When the ground truth label is 1, the model is expected to assign a high probability to this class. Similarly, when the ground truth label is 0, the model should assign a low probability to this class.
However, there is variation in the occurrence rates of different AUs, with some having low rates while others have high rates. Therefore, for AU detection, we usually use the weighted cross-entropy loss [11] ${L}_{wce}$, which is formulated as:
${L}_{wce}=-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}[y_{i}log\hat{y}_{i}+(1-{y}_{i})log(1-\hat{y}_{i})]$ (2)
where $w_{i}$=$\frac{(1/r_{i})n_{au}}{\sum_{i=1}^{n_{au}}(1/r_{i})}$ is the weight of the $i$-th AU determined by its occurrence frequency for balanced training. It permits reducing the impact of loss values associated with those AUs that are more commonly activated in the training set. Consequently, loss values arising from less frequently inactivating AUs are given greater significance during the training process. $r_{i}$ represents the occurrence frequency of the $i$-th AU, and $n_{au}$ is the total number of AUs.
The classification methods based on cross-entropy loss are generally applicable to most classification tasks. However, they suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization. This constraint becomes particularly problematic in the context of AU detection, where different AUs share a common encoder. For each AU of every sample, the predicted output $\hat{y}$ must satisfy the aforementioned condition, placing a significant burden on the fitting capability of the encoder. To address this challenge, we introduce a contrastive learning framework in the task of AU detection:
$\mathcal{L}_{wnce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{+}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (3)
$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))}]\}$
In the equation, $f(\cdot)$ serves as an encoder, mapping the features to a $d$-dimensional vector. The positive sample of $x$, denoted as $x^{+}$, corresponds to samples belonging to the same class as $x$. Typically, $x^{+}$ is obtained by applying various semantic-invariant operations such as random cropping or Gaussian blurring to $x$. On the other hand, $x^{-}$ represents negative samples of $x$, which are samples from different classes compared to $x$. $\text{sim}(f(x),f(x^{+}))$ measures the similarity between $x$ and its positive example $x^{+}$, while $\text{sim}(f(x),f(x^{-}_{j}))$ measures the similarity between the sample and its negative example. The term ${p}_{d}$ represents the data distribution, ${p}^{+}$ denotes the class conditional distribution of positive samples that have the same label as x, and ${p}^{-}$ represents the class conditional distribution of negative samples that have a different label from x. Inspired by Eq (2), we also add $w_{i}$ to Eq (3) to reduce the impact of loss values associated with those AUs that are more commonly activated in the training set.
The contrastive loss in Eq (3) distinguishes itself from the binary cross-entropy loss in Eq (1) in the following aspects: (1) The binary cross-entropy loss is computed based on the sample and its corresponding label y, requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the features of samples from different classes. This incentivizes the encoder to capture the differential information among different class samples, without the need to encode pixel-level information, enabling a lightweight encoder. (2) The binary cross-entropy loss constrains the absolute values of the predicted probabilities, while the contrastive loss constrains the difference in similarity between positive and negative sample pairs, without specifying the exact value. This is because the minimal value of $\mathcal{L}_{wnce}$ is 0, which is achieved when the difference between $\text{sim}((f(x),f(x^{+}))$ and $\text{sim}(f(x),f(x^{-}_{j}))$ approaches positive infinity for all $j\in{1,2,\cdots,N}$.
The contrastive learning encoder in our paper is shown in Fig. 2. Different image pairs are input into our backbone, which adopts Swin-transformer [50]. To ensure the fairness of the experimental comparison, The AUs Relationship-aware Node Feature Learning (ANFL) module the same as the encoder for the first stage training of baseline MEGraph [51]. Therefore, because of our discriminative contrastive learning paradigm, our encoder not only has half the parameters less than MEGraph, but also has better experimental results.
III-B Importance re-weighting strategy
To address the second challenge of the class imbalanced problem, we employ a importance re-weighting strategy to minority AUs. This helps to balance the representation of different AUs in the training data and improves the ability of models to accurately classify samples from all AUs. The improved loss ${L}_{aunce}$ is formulated as:
$\mathcal{L}_{aunce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{+}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (4)
$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))^{\beta}}]\}$
where $\beta\in(0,+\infty)$ is the importance weight. Without loss of generality, we consider the similarity calculation as $\text{sim}(f(x),f(x^{+}))=\exp(f(x)^{T}f(x^{+})/\tau)$, where $\tau$ is the temperature coefficient. The inclusion of importance weights transforms the gradient of similarity scores of positive pairs $\text{sim}(f(x),f(x^{+}))$ w.r.t $\mathcal{L}$ into the following form:
$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{+}))}=-\frac{1}{\tau}\sum_{j=1}^{N}P_{j}$ (5)
where $P_{j}=\frac{\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}{\exp(f(x)^{T}f(x^{+})/\tau)+\sum_{j=1}^{N}\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}$, and the gradient of similarity scores of negative pairs $\text{sim}(f(x),f(x^{-}))$ w.r.t $\mathcal{L}$ is
$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{-}))}=\frac{1}{\tau}P_{j}$ (6)
So, the relative gradient magnitudes of positive pairs and negative pairs can be expressed as follows:
$|\frac{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{+}))}{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{-}))}|=\frac{\sum_{j=1}^{N}\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}{\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}$ (7)
When $x$ belongs to the majority class, its positive sample is from the same majority class, while its negative sample belongs to the minority class. In this case, we aim to have smaller gradient magnitudes for the positive samples (majority class) to update fewer parameters, while the negative samples (minority class) should have larger gradient magnitudes to update more parameters. To achieve this, we begin to set $\beta\geq 1$, which increases the impact of parameter updates caused by minority class samples. On the other hand, when x belongs to the minority class, we begin to set $\beta\leq 1$ to reduce the influence of parameter updates caused by majority class samples. By adjusting the value of $\beta$, we can control the relative importance of positive and negative samples during the parameter update process, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. Theoretically, as long as $\beta_{minority}>\beta_{majority}$, the impact of the imbalanced problem can be mitigated.
III-C Enhanced Supervision Signal
Regarding the third challenge of AU labels tainted by noise and errors, we employ a combined approach that integrates label noisy learning and self-supervised learning techniques to obtain enhanced supervised signals for positive samples. This approach mitigates the adverse effects of noisy labels by introducing a controlled level of self-supervised signal and supervised signal, thereby enhancing the model’s robustness to variations and errors present in the labeled dataset.
III-C1 Supervised Signal
Our supervised signal comes from label noisy learning, which aims to address the issue of mislabeled or incorrectly annotated data by introducing mechanisms that account for label noise during training. One common technique is label smoothing, where a small amount of uncertainty is incorporated into the training process by replacing one-hot labels with smoothed probability distributions. Label smoothing allows the model to learn from the noisy labels while reducing the negative impact of label errors. However, it is challenging to directly apply label smoothing to contrastive learning as the computation of contrastive loss does not involve explicit labels. To overcome this challenge, we turn to the memory effect of neural networks. Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples. This memory effect is independent of the choice of neural network architecture. Inspired by this observation, for a given sample $x$ and M positive samples $\{x_{i}\}_{i=1}^{M}$ with noisy labels that belong to the same class, we compute the similarity scores between the representation of sample $x$ and the representations of all M positive samples. Based on these similarity scores, we select the positive feature representation with the highest similarity to sample $x$:
$f(x^{+})=f(x_{i}),\text{where}~{}i=\arg\max_{i}f(x)^{T}f(x_{i})$ (8)
By selecting the positive samples with the highest similarity, we prioritize clean and reliable positive samples while minimizing the influence of noisy positive samples during the training process.
III-C2 Self-supervised Signal
In addition to label noisy learning, we incorporate self-supervised learning techniques into the training process. By designing these auxiliary tasks carefully, we can encourage the model to learn more robust and invariant representations, which can help mitigate the effects of label noise and improve generalization performance. Specifically, our second method to obtain positive feature involves performing semantically invariant image augmentation on the anchor point x:
$f(x^{+})=f(\mathcal{T}(x))$ (9)
where $\mathcal{T}(\cdot)$ is a stochastic data augmentation module that applies random transformations to a given data example $x$, generating two correlated views referred to as $x_{1}^{+}$ and $x_{2}^{+}$, which we consider as a positive pair. Follow SimCLR [21], we employ a sequential application of three simple augmentations: random cropping and flipping followed by resizing back to the original size, random color distortions, and random Gaussian blur.
Our third method to obtain positive examples involving a mixture of positive feature representations, which can be seen as a mixup of positive samples to obtain the most representative positive instances.
$f(x^{+})=\frac{1}{M}\sum_{i=1}^{M}f(x_{i})$ (10)
The positive feature representation $f(x^{+})$ obtained from the equation above represents the class centroid of M positive examples $\{x_{i}\}_{i=1}^{M}$. In other words, it synthesizes a virtual representative positive sample, which is also referred to as feature augmentation. This method helps overcome the challenge of instance-level contrast, where each instance is treated as a separate class. By synthesizing a virtual sample that represents the class centroid, the feature representation of each instance is encouraged to be closer to the class centroid, which is benefit of capturing the essential characteristics of the positive class.
The aforementioned three methods of obtaining positive examples combine the use of label noisy learning for supervised signals and data augmentation for self-supervised signals. These three types of positive examples are randomly mixed with probabilities $p_{1}$, $p_{2}$, and $p_{3}$ respectively, where $p_{1}+p_{2}+p_{3}=1$. Our approach aims to improve the quality of supervised signals utilized in the training process. By integrating these techniques, we effectively mitigate the impact of AU labels tainted by noise and errors, increase the resilience of models to errors and variations in the labeled datasets, and ultimately enhance the generalization ability of models to unseen data. The pseudocode for the algorithm is provided below:
Input: Mini batch data $\mathcal{R}$, feature representation function $f(\cdot)$, $p_{1}$, $p_{2}$, and $p_{3}$, $\beta$
Output: Parameters $\Theta$.
for sample $x\in\mathcal{R}$ do
         Obtain positive sample set $\{x_{i}^{+}\}_{i=1}^{M}\in\mathcal{R}$;
         Obtain negative sample set $\{x_{j}^{-}\}_{j=1}^{N}\in\mathcal{R}$;
         p = random.uniform(0,1);
       if $p\leq p_{1}$ then
            Obtain $f(x^{+})$ via Eq (8)
        if $p_{1}<p\leq p_{1}+p_{2}$ then
            Obtain $f(x^{+})$ via Eq (9)
      else
            Obtain $f(x^{+})$ via Eq (10)
        Calculate AUNCE loss via Eq (4);
         Update $\Theta$.
Result: Parameters $\Theta$.
Algorithm 1 The proposed AUNCE algorithm
By addressing these three challenges, our proposed approach aims to improve the encoding capability of the shared encoder, enhance classification accuracy by addressing class imbalanced problem of each AU type, and enhance the robustness of models to AU labels tainted by noise and errors, ultimately leading to improved generalization performance in AU detection.
IV Evaluation
IV-A Experiment Setting
IV-A1 Datasets
In our experiments, we extensively utilize two well-established datasets for AU detection, namely BP4D [52] and DISFA [53]. Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations. For both datasets, we employ a three-fold cross-validation methodology for rigorous evaluation.
BP4D [52]: This dataset comprises 41 participants, including 23 females and 18 males. Each individual is featured in 8 sessions, recorded using both 2D and 3D videos. The dataset encompasses approximately 140,000 frames, each annotated with AU labels indicating occurrence or non-occurrence. Additionally, each frame is meticulously annotated with 68 landmarks detected using the SDM [54] method.
DISFA [53]: This dataset is composed of 27 videos, recorded from 15 men and 12 women. Each video comprises 4,845 frames, resulting in a total of 130,788 images. Frames in this dataset are annotated with AU intensities ranging from 0 to 5, and 68 landmarks are detected using the AAM [55] method. Frames with intensities equal to or greater than 2 are considered positive, while others are classified as negative.
IV-A2 Implementation Details
Initially, the 68 facial landmarks undergo a conversion into 49 facial internal landmarks, excluding those associated with the facial contour. Subsequently, the images undergo augmentation through similarity transformation, encompassing in-plane rotation, uniform scaling, and translation, utilizing the 49 facial landmarks. The resulting images are standardized to a resolution of 256$\times$256$\times$3 pixels and then randomly cropped into 224$\times$224$\times$3 pixels, with the addition of a random horizontal flip.
For the training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of $10^{-5}$ and 60 epochs. Due to our discriminative contrastive learning framework that sample pairs are selected in a same batch, so the batch size has a strong influence on the experimental results. The maximum batch size we can set is 32. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 8 for DISFA and 12 for BP4D respectively, the number K for choosing nearest neighbors in the ANFL is set to 4, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. V-C2.
In assessing the acquired representations, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [22, 56, 57]. This involves training a linear classifier atop the fixed base network, and the highest test accuracy across our models serves as a surrogate measure for the quality of the learned representations.
IV-A3 Evaluation Metrics
[…]
V Evaluation Metrics
TABLE I: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score.
Method Source AU Index Avg.
1 2 4 6 7 10 12 14 15 17 23 24
MAL [58] 2023 TAC 47.9 49.5 52.1 77.6 77.8 82.8 88.3 66.4 49.7 59.7 45.2 48.5 62.2
CLP [23] 2023 TIP 47.7 50.9 49.5 75.8 78.7 80.2 84.1 67.1 52.0 62.7 45.7 54.8 62.4
[…]
MMA-Net [41] 2023 PRL 52.5 50.9 58.3 76.3 75.7 83.8 87.9 63.8 48.7 61.7 46.5 54.4 63.4
HMP-PS [59] 2021 CVPR 53.1 46.1 56.0 76.5 76.9 82.1 86.4 64.8 51.5 63.0 49.9 54.5 63.4
AAR [60] 2023 TIP 53.2 47.7 56.7 75.9 79.1 82.9 88.6 60.5 51.5 61.9 51.0 56.8 63.8
SEV-Net [61] 2021 CVPR 58.2 50.4 58.3 81.9 73.9 87.7 87.5 61.6 52.6 62.2 44.6 47.6 63.9
SupHCL [62] 2022 ACM MM 52.8 45.7 61.6 79.5 79.3 84.7 86.9 67.6 51.4 62.5 48.6 52.3 64.4
KDRL [63] 2022 CVPR 53.3 47.4 56.2 79.4 80.7 85.1 89.0 67.4 55.9 61.9 48.5 49.0 64.5
AUNet [64] 2023 Proc.IEEE 58.0 48.2 62.4 76.4 77.5 83.4 88.5 63.3 52.0 65.5 52.1 52.3 65.0
MEGraph [51] 2022 IJCAI 52.7 44.3 60.9 79.9 80.1 85.3 89.2 69.4 55.4 64.4 49.8 55.1 65.5
AUNCE(Ours) - 54.3 49.5 60.8 78.3 78.1 84.4 89.1 66.6 56.3 64.8 49.6 56.5 65.7
TABLE II: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score.
Method Source AU Index Avg.
1 2 4 6 9 12 25 26
CLP [23] 2023 TIP 42.4 38.7 63.5 59.7 38.9 73.0 85.0 58.1 57.4
MAL [58] 2023 TAC 43.8 39.3 68.9 47.4 48.6 72.7 90.6 52.6 58.0
SEV-Net [61] 2021 CVPR 55.3 53.1 61.5 53.6 38.2 71.6 95.7 41.5 58.8
HMP-PS [59] 2021 CVPR 38.0 45.9 65.2 50.9 50.8 76.0 93.3 67.6 61.0
MEGraph [51] 2022 IJCAI 52.5 45.7 76.1 51.8 46.5 76.1 92.9 57.6 62.4
[…]
SupHCL [62] 2022 ACM MM 52.5 58.8 70.0 53.5 51.4 73.1 95.6 58.0 64.1
AAR [60] 2023 TIP 62.4 53.6 71.5 39.0 48.8 76.1 91.3 70.6 64.2
KDRL [63] 2022 CVPR 60.4 59.2 67.5 52.7 51.5 76.1 91.3 57.7 64.5
MMA-Net [41] 2023 PRL 63.8 54.8 73.6 39.2 61.5 73.1 92.3 70.5 66.0
AUNet [64] 2023 Proc.IEEE 60.3 59.1 69.8 48.4 53.0 79.7 93.5 64.7 66.1
AUNCE(Ours) - 61.8 58.9 74.9 49.7 56.2 73.5 92.1 64.2 66.4
V-A Comparison with State-of-the-Art Methods
While our discriminative contrastive learning primarily operates as a self-supervised framework, our experimental results on BP4D and DISFA still compare with some supervised frameworks, because we use AU labels assisting the process of selecting positive sample pairs. The focus of our comparison lies in evaluating state-of-the-art methods from the past three years, outlined as follows:
  • HMP-PS (CVPR 2021) [59] presents a pioneering approach to message passing neural networks incorporating performance-driven structures.
  • SEV-Net (CVPR 2021) [61] proposes a strategy for AU detection by leveraging semantic embedding and visual features.
[…]
  • SupHCL (ACM MM 2022) [62] introduces a method for AU detection that utilizes supervised hierarchical contrastive learning to ensure knowledge consistency across different facial images and AUs.
  • KDRL (CVPR 2022) [63] employs AU labeling rules from the Facial Action Coding System for self-supervised representation learning in AU detection.
  • MEGraph (IJCAI 2022) [51] focuses on deep learning of a unique graph to explicitly describe relationships between AUs in the target facial display.
  • MMA-Net (PRL 2023) [41] proposes a novel AU detection framework based on multiview mixed attention.
  • AUNet (Proc. IEEE 2023) [64] presents AU detection coupled with facial landmark detection, a simple yet robust baseline for landmark-based AU detection.
  • MAL (TAC 2023) [58] employs a meta learning approach to automatically select highly related facial expression samples using adaptative weights for training facial expression samples in AU detection.
  • AAR (TIP 2023) [60] introduces an adaptive attention regression network to regress the global attention map of each AU, guided by attention predefinition and AU detection.
  • CLP (TIP 2023) [23] aims to alleviate data scarcity issues by learning AU representations from a large pool of unlabeled facial videos in a contrastive learning paradigm.
The quantitative comparative results for the BP4D dataset are outlined in Table I. AUNCE exhibits superior performance in overall AU detection when contrasted with alternative approaches.
In comparison to MMA-Net, HMP-PS, AAR, SEV-Net, and KDRL, our approach exhibits a performance advantage of approximately 1.2%-2.3% in F1-score. This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing pertinent information for the training process. When contrasted with other contrastive learning methods such as CLP and SupHCL, our method demonstrates an average improvement of 3.3% and 1.3%, respectively. This improvement is attributed to our positive sample sampling strategy, which alleviates issues associated with AU labels tainted by noise and errors, enabling the encoder to extract differences among static image features more comprehensively.
[…]
Compared to the second-best approach, MEGraph, our method demonstrates superior performance by elevating the F1-score from 65.5% to 65.7%. It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly half. This highlights the applicability of AUNCE to more lightweight network structures and emphasizes that learning only the difference information among features is more potent than learning the pixel-level information of the entire image.
Comparable outcomes are evident in the DISFA dataset, as outlined in Table II, where AUNCE attains the highest overall performance in AU detection.
V-B Ablation Study
TABLE III: Quantitative results of ablation study on BP4D dataset.
Model BL PS IR F1-Score(%) Accuracy(%)
A $\checkmark$ 60.1 79.1
B $\checkmark$ $\checkmark$ 63.2(+3.1) 79.7(+0.6)
C $\checkmark$ $\checkmark$ 63.4(+3.3) 79.8(+0.7)
D $\checkmark$ $\checkmark$ $\checkmark$ 65.7(+5.6) 80.2(+1.1)
V-B1 Ablation Study
The quantitative results of our method with various component combinations are summarized in Table III, encompassing the baseline (BL) rooted in the new discriminative contrastive learning paradigm, utilizing images with the highest similarity to original images as positive sample pairs, the positive sample sampling strategy (PS), and the importance re-weighting strategy to minority AUs (IR). The merits of our proposed discriminative contrastive learning paradigm will be elaborated in Section V-B2.
Contribution of the positive sample sampling strategy: As outlined in Table III, the positive sample sampling strategy contributes to a 3.1% increase in F1-score and a 0.6% improvement in Accuracy compared to the baseline. This validates that selecting different types of positive sample pairs with appropriate probabilities can mitigate issues arising from AU labels tainted by noise and errors. The strategy achieves this by incorporating three types of positive sample pairs: the self-supervised signal pair, the highest similarity pair, and the mixed positive features pair. These pairs provide the models with diverse information from various views, thereby enhancing the generalization ability of the contrastive learning framework.
Effectiveness of the importance re-weighting strategy: The proposed importance re-weighting strategy to minority AUs results in a 3.3% increase in F1-score compared to model A. This illustrates that the strategy can change the backpropagation gradient of positive and negative samples, facilitating faster label category gradient updates with a smaller number of samples. Additionally, model D achieves an average increase of 2.5% in F1-score compared to model B, further supporting the efficacy of the importance re-weighting strategy.
Based on ablation studies, the introduction of each component proves beneficial for performance improvement, with IR emerging as the most effective component.
TABLE IV: Comparison of the weighted cross-entropy loss and AUNCE on BP4D and DISFA.
Dataset Loss MEGraph-stage1
F1-Score Accuracy
BP4D WCE 63.6 80.0
AUNCE 65.7(+2.1) 80.2(+0.2)
DISFA WCE 62.2 93.7
AUNCE 66.4(+4.2) 94.3(+0.6)
V-B2 Compared with weighted cross-entropy loss
As depicted in Table IV, AUNCE has exhibited superior performance compared to the traditional weighted cross-entropy loss (WCE) for AU detection, resulting in an increase in F1-score from 63.6% to 65.7% on BP4D and from 62.2% to 66.4% on DISFA. This highlights that capturing the distinctive features of different samples proves more effective than concentrating on pixel-level information. Consequently, our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance problem for each AU type, and bolsters model robustness against AU labels affected by noise and errors. This, in turn, leads to enhanced generalization performance in AU detection.
AUNCE can seamlessly integrate into various lightweight AU detection networks, serving as a replacement for generative models. The loss demonstrates robust compatibility in diverse model architectures.
V-C Hyperparameters Analysis
TABLE V: Quantitative results the probabilities of various positive sample pairs $p_{1},p_{2},p_{3},p_{4}$ on BP4D dataset.
$p_{1}$ $p_{2}$ $p_{3}$ $p_{4}$ F1-Score
Probabilities 1 0 0 0 60.1
0 1 0 0 42.0
0 0 1 0 63.2
0 0 0 1 17.6
0.1 0.1 0.8 0 63.8
0.2 0.2 0.6 0 65.7
0.3 0.3 0.4 0 64.3
0.4 0.4 0.2 0 58.8
V-C1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
In our paper, $p_{1}$ represents the probabilities of images with the highest similarity to original images, $p_{2}$ represents the probabilities of self-supervised images enhanced by simple augmentations, $p_{3}$ represents the probabilities of a mixture of all positive samples, and $p_{4}$ represents the probabilities of images with the lowest similarity to original images. Given that $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ address the problem of AU labels tainted by noise and errors, we systematically vary these probabilities from 0 to 1 to investigate their impact on our framework. As $p_{1}$ and $p_{2}$ are deemed reliable positive samples, we set their values to be the same during the variation.
The experimental results are presented in Table V. When $p_{1}$=$p_{2}$=0.2, $p_{3}$=0.6, and $p_{4}$=0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent errors and noise. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of error and noisy samples, thereby reinforcing the robustness of the training process.
Refer to caption (a)
Refer to caption (b)
Figure 3: Quantitative results of $\beta$ of different positive and negative samples for each AU type on BP4D dataset. From up to down: (a) $\beta$ for positive samples (the minority class), and (b) $\beta$ for negative samples (the majority class).
V-C2 Backpropagation rate controller $\beta$
As is well understood, the hyperparameter $\beta$ plays a crucial role in governing the relative importance of positive and negative samples during the parameter update process. It ensures that the model allocates more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. To explore its effects, we systematically vary the hyperparameter $\beta$ for the minority class from 0.8 to 1.8, with $\beta$ for the majority class fixed at 0.4. Additionally, we vary $\beta$ for the majority class from 0.2 to 1.2, keeping $\beta$ for the minority class fixed at 1.2.
The experimental results are depicted in Fig. 3. The optimal experimental outcome is observed when $\beta$ for the minority class is set to 1.2, and $\beta$ for the majority class is fixed at 0.4. Hence, when $\beta_{\text{minority}}>\beta_{\text{majority}}$, the impact of the imbalanced problem can be effectively mitigated.
VI Conclusion
In this paper, we introduce an innovative discriminative contrastive learning paradigm designed to replace traditional pixel-level learning in the domain of AU detection. Our proposed contrastive loss, referred to as AUNCE, shares similarities with InfoNCE, transforming the task of learning image data distribution information into the more effective learning of differentiated information among images. Additionally, we present a positive sample sampling strategy that introduces three types of positive sample pairs, addressing the challenge of AU labels tainted by noise and errors in two AU datasets. To further enhance the framework, we incorporate a importance re-weighting strategy specifically tailored to minority AUs. This strategy alters the backpropagation gradient of positive and negative samples, effectively addressing the imbalanced nature of each AU type. Experimental results underscore the superior performance of AUNCE in comparison to existing methods.
Acknowledgement
The computation is completed in the HPC Platform of Huazhong University of Science and Technology.

### v3

@@
 Abstract
-The predominant approach to facial action unit (AU) detection revolves around a supervised multi-label binary classification problem.
-Existing methodologies often encode pixel-level information of AUs, thereby imposing substantial demands on model complexity and expressiveness.
-Moreover, this practice elevates the susceptibility to overfitting due to the presence of noisy AU labels.
-In the present study, we introduce a contrastive learning framework enhanced by both supervised and self-supervised signals.
-The objective is to acquire discriminative features, deviating from the conventional pixel-level learning paradigm within the domain of AU detection.
-To address the challenge posed by noisy AU labels, we augment the supervised signal through the introduction of a self-supervised signal.
-This augmentation is achieved through positive sample sampling, encompassing three distinct types of positive sample pairs.
-Furthermore, to mitigate the imbalanced distribution of each AU type, we employ an importance re-weighting strategy tailored for minority AUs.
-The resulting loss, denoted as AUNCE, is proposed to encapsulate this strategy.
-Our experimental assessments, conducted on two widely-utilized benchmark datasets (BP4D and DISFA), underscore the superior performance of our approach compared to state-of-the-art methods in the realm of AU detection.
+Facial action unit (AU) detection has long encountered the challenge of detecting subtle feature differences when AUs activate.
+Existing methods often rely on encoding pixel-level information of AUs, which not only encodes additional redundant information but also leads to increased model complexity and limited generalizability.
+Additionally, the accuracy of AU detection is negatively impacted by the class imbalance issue of each AU type, and the presence of noisy and false AU labels.
+In this paper, we introduce a novel contrastive learning framework aimed for AU detection that incorporates both self-supervised and supervised signals, thereby enhancing the learning of discriminative features for accurate AU detection.
+To tackle the class imbalance issue, we employ a negative sample re-weighting strategy that adjusts the step size of updating parameters for minority and majority class samples.
+Moreover, to address the challenges posed by noisy and false AU labels, we employ a sampling technique that encompasses three distinct types of positive sample pairs.
+This enables us to inject self-supervised signals into the supervised signal, effectively mitigating the adverse effects of noisy labels.
+Our experimental assessments, conducted on four widely-utilized benchmark datasets (BP4D, DISFA, GFT and Aff-Wild2), underscore the superior performance of our approach compared to state-of-the-art methods of AU detection.
 Our code is available at https://github.com/Ziqiao-Shang/AUNCE.
@@
 I Introduction
-Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding.
-It has introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2], enabling their simulation.
-In contemporary times, AU detection [3, 4, 5, 6] has evolved into a technique utilizing computer vision algorithms to automatically identify and track AUs in videos or images.
-This advancement holds promise for diverse applications, including human-computer interaction [7], emotion analysis [8, 9], and monitoring in activities such as car driving [10].
-Refer to caption (a)
-Refer to caption (b)
-Refer to caption (c)
-Refer to caption (d)
-Figure 1: Illustration of our proposed positive sample sampling strategy.
-From left to right: (a) original images, (b) self-supervised images obtained by simple augmentations, (c) images with the highest similarity to (a), and (d) images with the lowest similarity to (a).
-We use (a) with (b), (c), mixture of all positive samples as our positive sample pairs.
-In recent years, a predominant trend in supervised frameworks for AU detection has been observed, as evidenced by various studies [11, 12, 13, 14, 15].
-These frameworks have predominantly employed the traditional binary cross-entropy loss, which serves as a metric for gauging the disparity between predicted and actual ground truth labels, providing a systematic means to guide the learning process.
-However, encoding the pixel-level information is notable for its heightened sensitivity to models characterized by a substantial number of parameters, as it endeavors to encode information at the pixel level within the input space.
-This imposes a notable demand on the fitting capability of the models.
-Additionally, pixel-level generative learning, while computationally intensive, may prove superfluous for AU detection tasks, given their nature as multi-label binary classification problems.
-The pursuit of pixel-level features not only introduces unwarranted model complexity but also augments the susceptibility to overfitting.
-Representation learning approaches rooted in the discriminative contrastive learning paradigm have demonstrated substantial promise, yielding state-of-the-art outcomes [16, 17, 18, 19, 20].
-The fundamental concept of contrastive learning involves maximizing the consistency of positive example features that are semantically similar while minimizing the consistency of negative example features that are semantically dissimilar [21].
-By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized, relieving the model from encoding pixel-level information [22].
-However, the direct integration of contrastive learning into AU detection tasks encounters two principal challenges.
-The first challenge pertains to the noisy label problem [23].
-Due to the weak amplitude of AU occurrence and the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit errors and noisy labels during the labeling process.
-Consequently, fitting models with noisy and erroneous samples significantly undermines the overall fitting ability of models.
-The second challenge is the class imbalance problems among different AUs [11].
-Most AU classes manifest a substantial imbalance between positive and negative samples, with inactivated AUs vastly outnumbering their activated counterparts.
-This imbalance not only diminishes the fitting efficiency of neural networks but also influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class.
-To tackle these challenges, we propose a simple contrastive learning framework to encode discriminative features between samples for AU detection, serving as a substitution for conventional pixel-level learning.
-This approach offers a pathway to lightweight model development within the domain of AU detection tasks.
-Moreover, in response to the issue of noisy labeling in AUs, we introduce self-supervised signals to augment the noise supervision signal.
-This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that positive examples derived from data augmentation with the self-supervised signal enhance robustness, as illustrated in Fig.
+Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2].
+AU detection is a multi-label binary classification problem.
+Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0.
+Nowadays, AU detection [3, 4, 5, 6, 7, 8, 9] has emerged as a computer vision technique to automatically identify AUs in videos or images.
+This advancement holds promise for diverse applications, including human-computer interaction [10], emotion analysis [11, 12], and car driving monitoring [13].
+In recent years, there has been a prevailing trend in supervised frameworks for AU detection [14, 15, 16, 17, 18, 19, 20].
+These studies have predominantly utilized the binary cross-entropy loss to strictly align predicted and ground truth labels, which is often unnecessary for classification tasks.
+This overemphasis on label alignment imposes excessive demands on the fitting capacity of models, consequently increasing the risk of overfitting.
+Moreover, these methods commonly focus on encoding entire pixel-level information of AUs, disregarding the fact that capturing subtle feature differences when AUs activate is sufficient.
+As a result, these learning methods place significant demands on the model complexity to encode excessive redundant information, ultimately limiting the generalization ability of the model.
+Figure 1: Overview of our proposed contrastive learning framework.
+Representation learning approaches rooted in contrastive learning have demonstrated substantial promise [21, 22, 23, 24, 25] for classification.
+This approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar [26].
+By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [27].
+However, the direct integration of contrastive learning into AU detection encounters two principal challenges.
+The first challenge is the class imbalance issue of each AU type [14].
+Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts.
+This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class.
+The second challenge pertains to the noisy and false label problem [28].
+Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
+To address these challenges, we propose a contrastive learning framework designed to encode discriminative features for AU detection, as depicted in Fig.
 1.
-Additionally, to mitigate the class imbalance problem among different types of AUs, we apply re-weighting to the contrastive loss.
-The resultant loss, resembling the commonly utilized InfoNCE, is termed AUNCE.
-Consequently, modifications are made to the gradient magnitude during back-propagation in the training process.
+This approach serves as a replacement for traditional pixel-level learning methods, which also offers a pathway to the lightweight model development within the field of AU detection.
+The resultant loss is termed AUNCE, resembling the commonly utilized InfoNCE.
+Moreover, to mitigate the class imbalance issue of each AU type, we apply a negative sample re-weighting strategy, making modifications to the gradient magnitude of minority and majority class samples during back-propagation by adding importance weight to the denominator of the resultant loss.
+Additionally, in response to the issue of noisy and false labeling in AUs, we introduce a positive sample samplying strategy to augment the supervision signal through self-supervised signals.
+This is achieved by establishing sampling ratios for various types of positive sample pairs, ensuring that more robust positive samples are derived from data augmentation with the self-supervised signal.
 The primary contributions of our work are succinctly summarized as follows:
-  • We propose a novel AUNCE loss to learn contrastive feature representations for AU detection.
-  • We introduce a strategy that incorporates self-supervised signal to address the issue of noisy AU labels.
-  • We employ an importance re-weighting strategy for minority AUs to tackle the class-imbalance problem.
-  • We validate the effectiveness of AUNCE by comparing it to state-of-the-art approaches for AU detection on two widely-used datasets, namely BP4D and DISFA.
+  • We propose a contrastive learning framework to learn discriminative feature representations for AU detection, and the resultant loss is named AUNCE.
+  • We employ a negative sample re-weighting strategy to tackle the class imbalance issue of each AU type.
+  • We introduce a positive sample samplying strategy to address the issue of noisy and false AU labels.
+  • We validate the effectiveness of AUNCE by comparing it to state-of-the-art approaches for AU detection on four widely-used datasets, namely BP4D, DISFA, GFT and Aff-Wild2.
 II Related Work
-We will address two key aspects closely linked to our work: deep contrastive learning and contrastive learning for AU detection.
+We provide a comprehensive overview of two areas closely connected to our research: deep contrastive learning and contrastive learning for AU detection.
 II-A Deep contrastive learning
@@
 This approach adheres to the learn-to-compare paradigm [31], wherein a distinction is made between observed data and noisy or erroneous data, preventing the model from engaging in the reconstruction of pixel-level information.
-Despite variations in representation encoders (denoted as $f$) and similarity measures across different tasks [27, 28], the underlying concept remains consistent: training the representation encoder $f$ by optimizing a contrastive loss [29], achieved through contrasting similar pairs ($x,x^{+}$) and dissimilar pairs ($x,x^{-}$).
 Prominent frameworks in this domain include SimCLR by Chen et al.
@@
 [32], employs a momentum network to maintain a queue of negative samples for efficient contrastive learning.
+Seyfi et al.
+[33] and Zhao et al.
+[34] have both proposed improved contrastive learning frameworks based on MoCo.
 Chen and He [35] mitigate undesired collapsing solutions, eschewing the use of negative sample pairs through a stop-gradient operation.
-Typically, these techniques are founded on the network architecture of quasi-twin neural networks.
-The image pairs utilized in the training process are derived from the same image, enlarged separately (positive sample pairs).
-The common practice involves obtaining positive samples through semantic-invariant operations on an anchor, coupled with heavy data augmentation techniques such as random cropping, flipping, image rotation, cutout, and color distortion.
-Conversely, image pairs are formed by enlarging different images separately (negative sample pairs).
-Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of numerous positive and negative sample pairs but also incorporates diverse losses for optimization.
+Despite variations and errors in representation encoders (denoted as $f$) and similarity measures across different tasks [36, 37], the underlying concept remains consistent: training the representation encoder $f$ by optimizing a contrastive loss [38], achieved through contrasting similar pairs ($x,x^{+}$) and dissimilar pairs ($x,x^{-}$).
+Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of positive and negative sample pairs but also incorporates diverse contrastive losses for optimization.
 For example, Oord et al.
@@
 [41] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
-As the technology of contrastive learning advances, there is a discernible shift in the construction of sample pairs for contrastive learning.
+There is a discernible shift in the construction of sample pairs for contrastive learning.
 In cases where data is labeled or partially labeled, positive samples can be constructed using images belonging to the same category, rather than being confined to identical images.
@@
 II-B Contrastive learning for AU detection
-While contrastive learning has garnered widespread popularity in recent years for AU detection, the predominant focus within AU detection methods has been on generative contrastive learning.
-Various approaches utilize a network with a parallel structure to directly compare features extracted from two or more sampled images.
-Wu et al.
-[35] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-affected facial images.
-Sun et al.
-[36] proposed an Emotion-aware Contrastive Learning framework designed to capture representations preserving sufficient AU-related information.
-Zhao et al.
-[37] devised a contrastive learning scheme aimed at creating a semantic-aware AU manifold, thereby reducing misalignment and lowering the dimensionality of facial action features.
+Contrastive learning was firstly designed as a pretext task [32] for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [42, 43, 44, 45, 46].
+For example, Niu et al.
+[42] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs).
 Liu et al.
 [45] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling.
-This multifaceted approach aimed to enhance the discriminative capabilities of AUs.
-Moreover, certain methodologies are specifically designed to incorporate contrastive losses tailored for AU detection tasks, aiming to discern differences among AU classes or different perspectives of the same AU class.
-Niu et al.
-[39] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs).
+Additionally, Chang et al.
+[46] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder.
+Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [47, 48, 49].
+for example, Wu et al.
+[47] introduced a contrastive feature learning method, enabling a Convolutional Neural Network (CNN) to acquire distinctive features that represent differences between neutral and AU-activated facial images.
+Chen et al.
+[49] introduced a supervised hierarchical contrastive learning method (SupHCL) aimed at enhancing AU detection by fostering knowledge consistency across distinct facial images and various AUs.
 In a similar vein, Suresh et al.
 [48] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
-Shang et al.
-[41] put forth a pixel-level multi-view contrastive loss, enhancing features through distinct cross-scale features.
-Additionally, Yang et al.
-[42] suggested a novel weighted multi-label contrastive learning loss tailored for AU detection.
-However, AU detection methods grounded in the discriminative contrastive learning paradigm have not found widespread adoption in recent research on AU detection, particularly in the context of unlabeled video sequences.
-Contemporary endeavors predominantly seek to generalize similarity learning between video frames within the discriminative contrastive learning paradigm, with the objective of acquiring meaningful video-level representations [43, 44, 45, 46, 47, 48].
-Nevertheless, these general self-supervised contrastive learning approaches generate representations that may not be specifically tailored to AU detection, leading to debates regarding the existence of a universal representation applicable to all downstream tasks [49].
-Despite the inherent temporal coherence and evolution characteristics present in unlabeled video sequences, Li et al.
-[23] have capitalized on these attributes by formulating a contrastive self-supervised learning method to encode discriminative AU representations from extensive amounts of unlabeled facial videos.
-However, the practical application of the discriminative contrastive learning paradigm for static images to AU detection frameworks remains theoretical, and the potential substitution of traditional generative cross-entropy loss by such learning methods is a subject of ongoing discussion.
+Different from the above static image-based approaches, some methods predominantly seek to generalize similarity learning between video frames within the discriminative paradigm [50, 51, 52, 53, 54, 55].
+One of the most representative methods is Li et al.
+[28] proposed, which have capitalized on these attributes by formulating a contrastive self-supervised learning method to encode discriminative AU representations from extensive amounts of unlabeled facial videos.
+In summary, contrastive learning has demonstrated promising results in various tasks.
+However, its direct application to the task of detecting subtle feature differences when AUs activate presents challenges: noisy and erroneous annotations, class imbalance in AUs, and difficulty in distinguishing hard negative samples.
 III Method
-In the Section.
-III, we mainly address three main challenges: 1.
-How to incentivize the encoder to capture differential information among AUs rather than pixel-level information, enabling a lightweight encoder through discriminative contrastive learning paradigm, instead of the traditional cross-entropy classification paradigm.
+In Section III, we present a contrastive learning framework specifically designed for the task of AU detection, aiming to incentivize the encoder to learn differential information between AUs from each sample pair, rather than focusing on pixel-level information.
+Subsequently, we confront two additional challenges: 1.
+Tackling the issue of class imbalance pertaining to each AU type in order to enhance the accuracy of our framework.
 2.
-How to solve the class imbalanced problem of each AU type to improve the accuracy of our framework.
-3.
-How to mitigate the problem of AU labels tainted by noise and errors, enhancing the robustness of our models to erroneously annotated data and thereby improving generalization performance of the models.
-Figure 2: Overview of the encoder of our contrastive learning framework.
-III-A Contrastive Learning
-To tackle the first challenge, we propose a contrastive learning framework that encourages the encoder to focus on capturing differences of each AU type instead of the pixel-level information of the whole face.
-By leveraging the relationships between samples and their corresponding AUs, the framework promotes the extraction of meaningful and discriminative features for AU detection.
-We investigate the feature representation of a specific AU, denoted as $x$.
-For AU detection which belongs to multi-label binary classification problem, the cross-entropy loss emerges as the primary choice for training models.
-This loss captures the discrepancy between predicted and ground truth labels, providing a principled approach to guide the learning process.
-Specifically, for a specific AU, the classification loss is calculated by comparing the features of samples with their corresponding labels.
-This enables the model to optimize its parameters with respect to the task of accurately classifying AUs.
-Mathematically, the typical cross-entropy loss can be expressed as follows:
-$\mathcal{L}_{\textsc{ce}}=y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})$ (1)
+Addressing the problem of noisy and false labels, aiming to bolster the robustness of our models against erroneously annotated data.
+III-A Contrastive learning framework
+For the task of AU detection, which falls under the category of multi-label binary classification, the prevalent approach for training models has been the utilization of cross-entropy loss.
+This loss function measures the disparity between predicted labels and ground truth labels.
+However, there exists variability in the occurrence rates of different AUs, with some AUs exhibiting low occurrence rates while others display high occurrence rates.
+Consequently, in contrast to the conventional cross-entropy loss, we employ a weighted cross-entropy loss [14] denoted as ${L}_{wce}$
+${L}_{wce}=-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}[y_{i}log\hat{y}_{i}+(1-{y}_{i})log(1-\hat{y}_{i})],$ (1)
 where $y_{i}$ denotes the ground truth value of occurrence (i.e., 0 or 1) of the $i$-th AU, and $\hat{y}_{i}$ is the corresponding predicted occurrence probability.
 This loss effectively guides the training process, encouraging the model to allocate higher probabilities to the correct class labels and lower probabilities to incorrect ones.
-In the case of sample $x_{i}$, the loss aims to minimize the discrepancy between predicted probabilities and the ground truth labels of the models.
-When the ground truth label is 1, the model is expected to assign a high probability to this class.
-Similarly, when the ground truth label is 0, the model should assign a low probability to this class.
-However, there is variation in the occurrence rates of different AUs, with some having low rates while others have high rates.
-Therefore, for AU detection, we usually use the weighted cross-entropy loss [11] ${L}_{wce}$, which is formulated as:
-${L}_{wce}=-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}[y_{i}log\hat{y}_{i}+(1-{y}_{i})log(1-\hat{y}_{i})]$ (2)
-where $w_{i}$=$\frac{(1/r_{i})n_{au}}{\sum_{i=1}^{n_{au}}(1/r_{i})}$ is the weight of the $i$-th AU determined by its occurrence frequency for balanced training.
+$w_{i}$=$\frac{(1/r_{i})n_{au}}{\sum_{i=1}^{n_{au}}(1/r_{i})}$ is the weight of the $i$-th AU determined by its occurrence frequency for balanced training.
 It permits reducing the impact of loss values associated with those AUs that are more commonly activated in the training set.
-Consequently, loss values arising from less frequently inactivating AUs are given greater significance during the training process.
 $r_{i}$ represents the occurrence frequency of the $i$-th AU, and $n_{au}$ is the total number of AUs.
-The classification methods based on cross-entropy loss are generally applicable to most classification tasks.
-However, they suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization.
+However, the classification methods based on cross-entropy loss suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization.
 This constraint becomes particularly problematic in the context of AU detection, where different AUs share a common encoder.
 For each AU of every sample, the predicted output $\hat{y}$ must satisfy the aforementioned condition, placing a significant burden on the fitting capability of the encoder.
-To address this challenge, we introduce a contrastive learning framework in the task of AU detection:
-$\mathcal{L}_{wnce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{+}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (3)
-$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))}]\}$
-In the equation, $f(\cdot)$ serves as an encoder, mapping the features to a $d$-dimensional vector.
+To address this challenge, we introduce a contrastive learning framework for AU detection.
+As shown in Fig.
+1.
+It comprises a positive sample mining module that integrates self-supervised signals with supervised signals, aiming to address the issue of noisy labels resulting from subtle differences in AUs.
+Additionally, a negative sample weighting module is included to tackle class imbalance and hard negative sample mining challenges.
+These modules will be discussed in detail in Sections III-B and III-C, respectively.
+The resultant loss is formulated as:
+$\mathcal{L}_{aunce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{+}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (2)
+$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\omega_{j}\cdot\text{sim}(f(x),f(x^{-}_{j}))}]\},$
+where $f(\cdot)$ serves as an encoder, mapping the features to a $d$ dimensional vector, represents the feature representation of a specific AU.
 The positive sample of $x$, denoted as $x^{+}$, corresponds to samples belonging to the same class as $x$.
-Typically, $x^{+}$ is obtained by applying various semantic-invariant operations such as random cropping or Gaussian blurring to $x$.
-On the other hand, $x^{-}$ represents negative samples of $x$, which are samples from different classes compared to $x$.
+Specifically, they are derived from three sources: samples from the same class, data augmentation of the anchor point, and feature augmentation of samples from the same class.
+$x^{-}$ represents negative samples of $x$, which are samples from different classes compared to $x$.
 $\text{sim}(f(x),f(x^{+}))$ measures the similarity between $x$ and its positive example $x^{+}$, while $\text{sim}(f(x),f(x^{-}_{j}))$ measures the similarity between the sample and its negative example.
-The term ${p}_{d}$ represents the data distribution, ${p}^{+}$ denotes the class conditional distribution of positive samples that have the same label as x, and ${p}^{-}$ represents the class conditional distribution of negative samples that have a different label from x.
-Inspired by Eq (2), we also add $w_{i}$ to Eq (3) to reduce the impact of loss values associated with those AUs that are more commonly activated in the training set.
-The contrastive loss in Eq (3) distinguishes itself from the binary cross-entropy loss in Eq (1) in the following aspects: (1) The binary cross-entropy loss is computed based on the sample and its corresponding label y, requiring the encoding of pixel-level information in order to align with the label.
+The term ${p}_{d}$ represents the data distribution, ${p}^{+}$ and ${p}^{-}$ respectively denote the class conditional distribution of positive and negative samples of $x$.
+Inspired by Eq (1), we introduce $w_{i}$ in Eq (2) to mitigate the influence of loss values related to commonly activated AUs in the training set.
+Additionally, we incorporate $\omega_{j}$ to re-weight the significance of each negative sample.
+The contrastive loss ${L}_{aunce}$ distinguishes itself from weighted cross-entropy loss ${L}_{wce}$ in the following aspects: 1.
+The weighted cross-entropy loss is computed based on the sample and its corresponding label $y$, requiring the encoding of pixel-level information in order to align with the label.
 In contrast, the contrastive loss does not involve explicit labels but is computed based on the features of samples from different classes.
 This incentivizes the encoder to capture the differential information among different class samples, without the need to encode pixel-level information, enabling a lightweight encoder.
-(2) The binary cross-entropy loss constrains the absolute values of the predicted probabilities, while the contrastive loss constrains the difference in similarity between positive and negative sample pairs, without specifying the exact value.
-This is because the minimal value of $\mathcal{L}_{wnce}$ is 0, which is achieved when the difference between $\text{sim}((f(x),f(x^{+}))$ and $\text{sim}(f(x),f(x^{-}_{j}))$ approaches positive infinity for all $j\in{1,2,\cdots,N}$.
-The contrastive learning encoder in our paper is shown in Fig.
 2.
-Different image pairs are input into our backbone, which adopts Swin-transformer [50].
-To ensure the fairness of the experimental comparison, The AUs Relationship-aware Node Feature Learning (ANFL) module the same as the encoder for the first stage training of baseline MEGraph [51].
-Therefore, because of our discriminative contrastive learning paradigm, our encoder not only has half the parameters less than MEGraph, but also has better experimental results.
-III-B Importance re-weighting strategy
-To address the second challenge of the class imbalanced problem, we employ a importance re-weighting strategy to minority AUs.
-This helps to balance the representation of different AUs in the training data and improves the ability of models to accurately classify samples from all AUs.
-The improved loss ${L}_{aunce}$ is formulated as:
-$\mathcal{L}_{aunce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{+}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (4)
-$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))^{\beta}}]\}$
-where $\beta\in(0,+\infty)$ is the importance weight.
+The weighted cross-entropy loss imposes constraints on the absolute values of the predicted probabilities.
+However, for classification tasks, this is often unnecessary and can lead to overfitting as well as demanding higher fitting capability from the encoder.
+The contrastive loss focuses on constraining the similarity difference between positive and negative sample pairs without explicitly specifying their exact values.
+This is because the minimal value of $\mathcal{L}_{aunce}$ is 0, which is achieved when $\text{sim}((f(x),f(x^{+}))-\text{sim}(f(x),f(x^{-}_{j}))\rightarrow+\infty$ for all negative sample index $j\in{1,2,\cdots,N}$.
+To ensure a fair comparison with conventional pixel-level learning paradigms, we utilize the network structure from the first training stage of MEGraph [56] as our contrastive learning encoder, as illustrated in Fig.
+2.
+The input to the encoder consists of an contrastive larning image pair consisting of an anchor and a random sample, and the encoder is optimized by AUNCE only.
+Our backbone adopts Swin-transformer [57], and the AUs Relationship-aware Node Feature Learning (ANFL) module is composed of two integral components: the AU-specific Feature Generator (AFG) and the Facial Graph Generator (FGG).
+The AFG independently generates a unique representation for each AU.
+Building on these representations, the FGG constructs an optimal graph for each facial display to accurately recognize all target AUs.
+To accomplish this, the FGG ensures that the AFG encodes task-specific associations among AUs into their respective representations.
+Figure 2: Overview of the encoder of our contrastive learning framework.
+III-B Negative sample re-weighting strategy
+To address the second challenge of the class imbalance issue, we employ a negative sample re-weighting strategy to minority AUs.
+This also contributes to hard negative sample mining.
+Inspired by HCL [58], we apply the following weighting scheme to each negative sample $x_{j}^{-}$:
+$\omega_{j}=\text{sim}(f(x),f(x^{-}_{j}))^{\beta},$ (3)
+where $\beta\in(0,+\infty)$ can be viewed as the importance weight if we assume that the negative samples follow a von Mises-Fisher (vMF) distribution. The resulting AUNCE loss is:
+$\mathcal{L}_{aunce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{-}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (4)
+$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))^{\beta+1}}]\}.$
+Next, we analyze how this weighting scheme addresses the class imbalance issue and facilitates the mining of hard negative samples.
 Without loss of generality, we consider the similarity calculation as $\text{sim}(f(x),f(x^{+}))=\exp(f(x)^{T}f(x^{+})/\tau)$, where $\tau$ is the temperature coefficient.
-The inclusion of importance weights transforms the gradient of similarity scores of positive pairs $\text{sim}(f(x),f(x^{+}))$ w.r.t $\mathcal{L}$ into the following form:
-$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{+}))}=-\frac{1}{\tau}\sum_{j=1}^{N}P_{j}$ (5)
+The inclusion of importance weights transforms the gradient of $\mathcal{L}$ w.r.t similarity scores of positive pairs into the following form:
+$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{+}))}=-\frac{1}{\tau}\sum_{j=1}^{N}P_{j},$ (5)
 where $P_{j}=\frac{\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}{\exp(f(x)^{T}f(x^{+})/\tau)+\sum_{j=1}^{N}\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}$, and the gradient of similarity scores of negative pairs $\text{sim}(f(x),f(x^{-}))$ w.r.t $\mathcal{L}$ is
-$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{-}))}=\frac{1}{\tau}P_{j}$ (6)
-So, the relative gradient magnitudes of positive pairs and negative pairs can be expressed as follows:
-$|\frac{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{+}))}{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{-}))}|=\frac{\sum_{j=1}^{N}\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}{\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}$ (7)
-When $x$ belongs to the majority class, its positive sample is from the same majority class, while its negative sample belongs to the minority class.
+$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{-}))}=\frac{1}{\tau}P_{j}.$ (6)
+Therefore, the ratio between the gradient magnitude of negative examples and positive examples is given by:
+$|\frac{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{-}))}{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{+}))}|=\frac{\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}{\sum_{j=1}^{N}\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}.$ (7)
+Above equation follows the Boltzmann distribution.
+As the value of $\beta$ increases, the entropy of the distribution decreases significantly [59], leading to a sharper distribution in the region of high similarity.
+This has two consequences: Firstly, it imposes larger penalties on hard negative samples that are in close proximity to the anchor data point, enabling the model to learn more precise decision boundaries from subtle feature differences when AUs activate.
+Secondly, it controls the relative contribution of negative samples to the parameter updates, providing a means to address class imbalance issues.
+Specifically, when anchor data point $x$ belongs to the majority class, its positive sample belongs to the majority class, while its negative sample belongs to the minority class.
 In this case, we aim to have smaller gradient magnitudes for the positive samples (majority class) to update fewer parameters, while the negative samples (minority class) should have larger gradient magnitudes to update more parameters.
-To achieve this, we begin to set $\beta\geq 1$, which increases the impact of parameter updates caused by minority class samples.
-On the other hand, when x belongs to the minority class, we begin to set $\beta\leq 1$ to reduce the influence of parameter updates caused by majority class samples.
-By adjusting the value of $\beta$, we can control the relative importance of positive and negative samples during the parameter update process, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly.
-Theoretically, as long as $\beta_{minority}>\beta_{majority}$, the impact of the imbalanced problem can be mitigated.
-III-C Enhanced Supervision Signal
+To achieve this, we begin to set $\beta\geq 1$, which increases the impact of parameter updates caused by negative samples (minority class).
+On the other hand, when $x$ belongs to minority class, we begin to set $\beta\leq 1$ to reduce the influence of parameter updates caused by majority class samples.
+By adjusting the value of $\beta$, we can control the relative importance of positive and negative samples during the parameter updating, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly.
+This helps to balance the representation of different AUs in training set and improves the model ability to precisely classify all AU samples.
+Figure 3: Illustration of our proposed positive sample samplying strategy.
+III-C Positive sample samplying strategy
 Regarding the third challenge of AU labels tainted by noise and errors, we employ a combined approach that integrates label noisy learning and self-supervised learning techniques to obtain enhanced supervised signals for positive samples.
-This approach mitigates the adverse effects of noisy labels by introducing a controlled level of self-supervised signal and supervised signal, thereby enhancing the model’s robustness to variations and errors present in the labeled dataset.
+As shown in Fig.
+3, this approach mitigates the adverse effects of noisy labels by simultaneously introducing a controlled level of self-supervised signal and supervised signal, thereby enhancing the robustness of model to variations and errors present in the labeled dataset.
 III-C1 Supervised Signal
-Our supervised signal comes from label noisy learning, which aims to address the issue of mislabeled or incorrectly annotated data by introducing mechanisms that account for label noise during training.
+Our supervised signal comes from label noisy learning, which aims to address the issue of mislabeled data by introducing mechanisms that account for label noise during training.
 One common technique is label smoothing, where a small amount of uncertainty is incorporated into the training process by replacing one-hot labels with smoothed probability distributions.
@@
 To overcome this challenge, we turn to the memory effect of neural networks.
-Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples.
+Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples [60, 61, 62].
 This memory effect is independent of the choice of neural network architecture.
-Inspired by this observation, for a given sample $x$ and M positive samples $\{x_{i}\}_{i=1}^{M}$ with noisy labels that belong to the same class, we compute the similarity scores between the representation of sample $x$ and the representations of all M positive samples.
-Based on these similarity scores, we select the positive feature representation with the highest similarity to sample $x$:
-$f(x^{+})=f(x_{i}),\text{where}~{}i=\arg\max_{i}f(x)^{T}f(x_{i})$ (8)
+The positive samples that are easier fit are those that are closer to the anchor point and have higher similarity.
+Inspired by this observation, for a given sample $x$ and M positive samples ${x}_{i=1}^{M}$ with noisy labels belonging to the same class, we calculate the similarity scores between the representation of sample $x$ and the representations of all M positive samples.
+We then select the positive feature representation with the highest similarity to sample $x$:
+$f(x_{h}^{+})=f(x_{i}),\text{where}~{}i=\arg\max_{i}f(x)^{T}f(x_{i}).$ (8)
 By selecting the positive samples with the highest similarity, we prioritize clean and reliable positive samples while minimizing the influence of noisy positive samples during the training process.
@@
 Specifically, our second method to obtain positive feature involves performing semantically invariant image augmentation on the anchor point $x$:
-$f(x^{+})=f(\mathcal{T}(x))$ (9)
-where $\mathcal{T}(\cdot)$ is a stochastic data augmentation module that applies random transformations to a given data example $x$, generating two correlated views referred to as $x_{1}^{+}$ and $x_{2}^{+}$, which we consider as a positive pair.
+$f(x_{s}^{+})=f(\mathcal{T}(x)),$ (9)
+where $\mathcal{T}(\cdot)$ is a stochastic data augmentation module that applies random transformations to a given data example $x$, generating two correlated views referred to as $x_{s1}^{+}$ and $x_{s2}^{+}$, which we consider as a positive pair.
 Follow SimCLR [22], we employ a sequential application of three simple augmentations: random cropping and flipping followed by resizing back to the original size, random color distortions, and random Gaussian blur.
-Our third method to obtain positive examples involving a mixture of positive feature representations, which can be seen as a mixup of positive samples to obtain the most representative positive instances.
-$f(x^{+})=\frac{1}{M}\sum_{i=1}^{M}f(x_{i})$ (10)
-The positive feature representation $f(x^{+})$ obtained from the equation above represents the class centroid of M positive examples $\{x_{i}\}_{i=1}^{M}$.
+Our third method to obtain positive examples involving a mixture of positive feature representations, which can be seen as a mixup of positive samples to obtain the most representative positive instances:
+$f(x_{m}^{+})=\frac{1}{M}\sum_{i=1}^{M}f(x_{i}).$ (10)
+The positive feature representation $f(x_{m}^{+})$ obtained from the equation above represents the class centroid of M positive examples $\{x\}_{i=1}^{M}$.
 In other words, it synthesizes a virtual representative positive sample, which is also referred to as feature augmentation.
@@
 Our approach aims to improve the quality of supervised signals utilized in the training process.
-By integrating these techniques, we effectively mitigate the impact of AU labels tainted by noise and errors, increase the resilience of models to errors and variations in the labeled datasets, and ultimately enhance the generalization ability of models to unseen data.
-The pseudocode for the algorithm is provided below:
+By integrating these techniques, we effectively mitigate the impact of noisy and false labels, increase the resilience of models to errors and variations in labeled datasets, and ultimately enhance the generalization ability of models to unseen data.
+By addressing the above three challenges, our approach improves the encoding capability of the shared encode, ultimately leading to improved generalization performance in AU detection.
+The pseudocode for the algorithm is provided in Algorithm.
+1.
 Input: Mini batch data $\mathcal{R}$, feature representation function $f(\cdot)$, $p_{1}$, $p_{2}$, and $p_{3}$, $\beta$
@@
        if $p\leq p_{1}$ then
-            Obtain $f(x^{+})$ via Eq (8)
+            Obtain $f(x_{h}^{+})$ via Eq (8)
         if $p_{1}<p\leq p_{1}+p_{2}$ then
-            Obtain $f(x^{+})$ via Eq (9)
+            Obtain $f(x_{s}^{+})$ via Eq (9)
       else
-            Obtain $f(x^{+})$ via Eq (10)
+            Obtain $f(x_{m}^{+})$ via Eq (10)
         Calculate AUNCE loss via Eq (4);
@@
 Algorithm 1 The proposed AUNCE algorithm
-By addressing these three challenges, our proposed approach aims to improve the encoding capability of the shared encoder, enhance classification accuracy by addressing class imbalanced problem of each AU type, and enhance the robustness of models to AU labels tainted by noise and errors, ultimately leading to improved generalization performance in AU detection.
 IV Evaluation
 IV-A Experiment Setting
+TABLE I: AU occurrence rates (%) in the training sets of different datasets. ”-” denotes the AU is not annotated in this dataset.
+Dataset AU Index
+1 2 4 6 7 9 10 12 14 15 17 23 24 25 26
+BP4D [63] 21.1 17.1 20.3 46.2 54.9 - 59.4 56.2 46.6 16.9 34.4 16.5 15.2 - -
+DISFA [64] 5.0 4.0 15.0 8.1 - 4.3 - 13.2 - - - - - 27.8 8.9
+GFT [65] 3.7 13.5 3.7 28.3 - - 24.6 29.3 3.1 10.7 - 25.0 14.1 - -
+Aff-Wild2 [66] 11.9 5.1 16.0 26.5 39.9 - 34.5 24.3 - 2.8 - 3.1 2.8 62.8 7.6
 IV-A1 Datasets
-In our experiments, we extensively utilize two well-established datasets for AU detection, namely BP4D [52] and DISFA [53].
+In our experiments, we extensively utilize three well-established datasets in constrained scenarios for AU detection, namely BP4D [63], DISFA [64], GFT [65], and one well-established dataset in unconstrained scenarios, namely Aff-Wild2 [66].
 Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations.
-For both datasets, we employ a three-fold cross-validation methodology for rigorous evaluation.
 BP4D [63]: This dataset comprises 41 participants, including 23 females and 18 males.
@@
 The dataset encompasses approximately 140,000 frames, each annotated with AU labels indicating occurrence or non-occurrence.
-Additionally, each frame is meticulously annotated with 68 landmarks detected using the SDM [54] method.
+Additionally, each frame is meticulously annotated with 68 facial landmarks detected using the SDM [67] method.
+For BP4D dataset, we employ a three-fold cross-validation methodology for rigorous evaluation.
 DISFA [64]: This dataset is composed of 27 videos, recorded from 15 men and 12 women.
 Each video comprises 4,845 frames, resulting in a total of 130,788 images.
-Frames in this dataset are annotated with AU intensities ranging from 0 to 5, and 68 landmarks are detected using the AAM [55] method.
+Frames in this dataset are annotated with AU intensities ranging from 0 to 5, and 68 facial landmarks are detected using the AAM [68] method.
 Frames with intensities equal to or greater than 2 are considered positive, while others are classified as negative.
+We also employ the same three-fold cross-validation methodology as BP4D dataset for rigorous evaluation of DISFA dataset.
+GFT [65]: This dataset comprises 96 participants, organized into 32 groups of three individuals, engaged in unscripted conversations.
+Each participant is recorded via video, with the majority of frames capturing moderate out-of-plane poses.
+Annotations for each frame include 10 AUs, alongside 68 facial landmarks identified using the ZFace [69] tool.
+In accordance with the standard training and testing partitions, our training set consists of 78 participants, amounting to approximately 108,000 frames, while the testing set includes 18 participants, totaling around 24,600 frames.
+Aff-Wild2 [66]: This dataset is an extensive in-the-wild facial expression dataset, with videos sourced from YouTube, encompassing a broad range of ages, ethnicities, professions, emotions, poses, lighting conditions, and occlusions.
+This dataset comprises 305 videos, totaling approximately 1,390,000 frames for the training set, and 105 videos, with around 440,000 frames for the validation set.
+Each frame is annotated with 12 AUs.
+For facial landmark annotations, we utilize the Dlib library [70], [71] for detecting 68 facial landmarks per frame.
+Following the methodology of Zhang et al.
+[72], we train our AUNCE on the training set and evaluate it on the validation set.
+The occurrence rates of AUs in the training sets of these datasets are presented in Table I.
+It is evident from the table that the class imbalance issue of each type of AU is prevalent across all datasets, especially on the DISFA, GFT and Aff-Wild2 datasets.
 IV-A2 Implementation Details
@@
 Due to our discriminative contrastive learning framework that sample pairs are selected in a same batch, so the batch size has a strong influence on the experimental results.
-The maximum batch size we can set is 32.
-In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 8 for DISFA and 12 for BP4D respectively, the number K for choosing nearest neighbors in the ANFL is set to 4, and the temperature coefficient $\tau$ is 0.5.
+In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D and 8 for DISFA respectively, the number K for choosing nearest neighbors in the ANFL [56] is set to 4, and the temperature coefficient $\tau$ is 0.5.
 The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section.
-V-C2.
-In assessing the acquired representations, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [22, 56, 57].
+IV-D2.
+In assessing the acquired representations, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [27, 73, 74, 22].
 This involves training a linear classifier atop the fixed base network, and the highest test accuracy across our models serves as a surrogate measure for the quality of the learned representations.
@@
 […]
-V Evaluation Metrics
-TABLE I: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score.
+TABLE II: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score(%).
 Method Source AU Index Avg.
 1 2 4 6 7 10 12 14 15 17 23 24
+EAC-Net [7] 2018 TPAMI 39.0 35.2 48.6 76.1 72.9 81.9 86.2 58.8 37.5 59.1 35.9 35.8 55.9
+ROI [4] 2017 CVPR 36.2 31.6 43.4 77.1 73.7 85.0 87.0 62.6 45.7 58.0 38.3 37.4 56.4
+ARL [15] 2019 TAC 45.8 39.8 55.1 75.7 77.2 82.3 86.6 58.8 47.6 62.1 47.4 55.4 61.1
 MAL [75] 2023 TAC 47.9 49.5 52.1 77.6 77.8 82.8 88.3 66.4 49.7 59.7 45.2 48.5 62.2
-CLP [23] 2023 TIP 47.7 50.9 49.5 75.8 78.7 80.2 84.1 67.1 52.0 62.7 45.7 54.8 62.4
 […]
+GLEE-Net [20] 2024 TCSVT 60.6 44.4 61.0 80.6 78.7 85.4 88.1 64.9 53.7 65.1 47.7 58.5 65.7
 MMA-Net [76] 2023 PRL 52.5 50.9 58.3 76.3 75.7 83.8 87.9 63.8 48.7 61.7 46.5 54.4 63.4
+SEV-Net [77] 2021 CVPR 58.2 50.4 58.3 81.9 73.9 87.7 87.5 61.6 52.6 62.2 44.6 47.6 63.9
 HMP-PS [78] 2021 CVPR 53.1 46.1 56.0 76.5 76.9 82.1 86.4 64.8 51.5 63.0 49.9 54.5 63.4
 AAR [79] 2023 TIP 53.2 47.7 56.7 75.9 79.1 82.9 88.6 60.5 51.5 61.9 51.0 56.8 63.8
-SEV-Net [61] 2021 CVPR 58.2 50.4 58.3 81.9 73.9 87.7 87.5 61.6 52.6 62.2 44.6 47.6 63.9
+ISTR [80] 2021 ACM MM 54.0 46.0 55.7 79.4 78.8 84.5 87.0 67.0 55.6 63.1 50.7 55.3 64.8
+MEGraph [56] 2022 IJCAI 52.7 44.3 60.9 79.9 80.1 85.3 89.2 69.4 55.4 64.4 49.8 55.1 65.5
+SimCLR [22] 2020 ICML 38.0 36.4 37.2 66.6 64.7 76.2 76.2 51.1 29.8 56.1 27.5 37.7 49.8
+MoCo [32] 2020 CVPR 30.8 41.3 42.1 70.2 70.4 78.7 82.5 53.3 25.2 59.1 31.5 34.3 51.6
+CLP [28] 2023 TIP 47.7 50.9 49.5 75.8 78.7 80.2 84.1 67.1 52.0 62.7 45.7 54.8 62.4
+BTFM [81] 2023 CVPR 57.4 52.6 64.6 79.3 81.5 82.7 85.6 67.9 47.3 58.0 47.0 44.9 64.1
 SupHCL [49] 2022 ACM MM 52.8 45.7 61.6 79.5 79.3 84.7 86.9 67.6 51.4 62.5 48.6 52.3 64.4
-KDRL [63] 2022 CVPR 53.3 47.4 56.2 79.4 80.7 85.1 89.0 67.4 55.9 61.9 48.5 49.0 64.5
-AUNet [64] 2023 Proc.IEEE 58.0 48.2 62.4 76.4 77.5 83.4 88.5 63.3 52.0 65.5 52.1 52.3 65.0
-MEGraph [51] 2022 IJCAI 52.7 44.3 60.9 79.9 80.1 85.3 89.2 69.4 55.4 64.4 49.8 55.1 65.5
-AUNCE(Ours) - 54.3 49.5 60.8 78.3 78.1 84.4 89.1 66.6 56.3 64.8 49.6 56.5 65.7
-TABLE II: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score.
+KSRL [46] 2022 CVPR 53.3 47.4 56.2 79.4 80.7 85.1 89.0 67.4 55.9 61.9 48.5 49.0 64.5
+CLEF [82] 2023 ICCV 55.8 46.8 63.3 79.5 77.6 83.6 87.8 67.3 55.2 63.5 53.0 57.8 65.9
+AUNCE(Ours) - 53.6 49.8 61.6 78.4 78.8 84.7 89.6 67.4 55.1 65.4 50.9 58.0 66.1
+(±$\pm 0.15$) (±$\pm 0.12$) (±$\pm 0.17$) (±$\pm 0.21$) (±$\pm 0.19$) (±$\pm 0.22$) (±$\pm 0.22$) (±$\pm 0.17$) (±$\pm 0.14$) (±$\pm 0.15$) (±$\pm 0.14$) (±$\pm 0.15$) (±$\pm 0.17$)
+TABLE III: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score(%).
 Method Source AU Index Avg.
 1 2 4 6 9 12 25 26
+EAC-Net [7] 2018 TPAMI 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
+ROI [4] 2017 CVPR 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
+ARL [15] 2019 TAC 43.9 42.1 63.6 41.8 40.0 76.2 95.2 66.8 58.7
+MAL [75] 2023 TAC 43.8 39.3 68.9 47.4 48.6 72.7 90.6 52.6 58.0
+[…]
+GLEE-Net [20] 2024 TCSVT 61.9 54.0 75.8 45.9 55.7 77.6 92.9 60.0 65.5
+SEV-Net [77] 2021 CVPR 55.3 53.1 61.5 53.6 38.2 71.6 95.7 41.5 58.8
+MMA-Net [76] 2023 PRL 63.8 54.8 73.6 39.2 61.5 73.1 92.3 70.5 66.0
+HMP-PS [78] 2021 CVPR 38.0 45.9 65.2 50.9 50.8 76.0 93.3 67.6 61.0
+ISTR [80] 2021 ACM MM 47.5 53.3 64.4 51.8 44.4 74.7 92.1 60.7 61.1
+MEGraph [56] 2022 IJCAI 52.5 45.7 76.1 51.8 46.5 76.1 92.9 57.6 62.4
+AAR [79] 2023 TIP 62.4 53.6 71.5 39.0 48.8 76.1 91.3 70.6 64.2
+SIMCLR [22] 2020 ICML 21.2 23.3 47.5 42.4 35.5 66.8 81.5 52.7 46.4
+MoCo [32] 2020 CVPR 22.7 18.2 45.9 45.4 34.1 72.9 83.4 54.5 47.1
 CLP [28] 2023 TIP 42.4 38.7 63.5 59.7 38.9 73.0 85.0 58.1 57.4
-MAL [58] 2023 TAC 43.8 39.3 68.9 47.4 48.6 72.7 90.6 52.6 58.0
-SEV-Net [61] 2021 CVPR 55.3 53.1 61.5 53.6 38.2 71.6 95.7 41.5 58.8
-HMP-PS [59] 2021 CVPR 38.0 45.9 65.2 50.9 50.8 76.0 93.3 67.6 61.0
-MEGraph [51] 2022 IJCAI 52.5 45.7 76.1 51.8 46.5 76.1 92.9 57.6 62.4
+BTFM [81] 2023 CVPR 41.5 44.9 60.3 51.5 50.3 70.4 91.3 55.3 58.2
+SupHCL [49] 2022 ACM MM 52.5 58.8 70.0 53.5 51.4 73.1 95.6 58.0 64.1
+KSRL [46] 2022 CVPR 60.4 59.2 67.5 52.7 51.5 76.1 91.3 57.7 64.5
+CLEF [82] 2023 ICCV 64.3 61.8 68.4 49.0 55.2 72.9 89.9 57.0 64.8
+AUNCE(Ours) - 61.8 58.9 74.9 49.7 56.2 73.5 92.1 64.2 66.4
+(±$\pm 0.37$) (±$\pm 0.22$) (±$\pm 0.42$) (±$\pm 0.31$) (±$\pm 0.36$) (±$\pm 0.49$) (±$\pm 0.42$) (±$\pm 0.28$) (±$\pm 0.36$)
+TABLE IV: Quantitative comparison results on GFT dataset, where the evaluation metric is F1-score(%).
+Method Source AU Index Avg.
+1 2 4 6 10 12 14 15 23 24
+AlexNet [83] 2012 NIPS 44 46 2 73 72 82 5 19 43 42 42.8
+LSVM [84] 2008 JMLR 38 32 13 67 64 78 15 29 49 44 42.9
+TCAE [85] 2019 CVPR 43.9 49.5 6.3 71.0 76.2 79.5 10.7 28.5 34.5 41.7 44.2
+CPAM [86] 2020 TBIOM 43.7 44.9 19.8 74.6 76.5 79.8 50.0 33.9 16.8 12.9 45.3
+TAE [87] 2022 TPAMI 46.3 48.8 13.4 76.7 74.8 81.8 19.9 42.3 50.6 50.0 50.5
+EAC-Net [7] 2018 TPAMI 15.5 56.6 0.1 81.0 76.1 84.0 0.1 38.5 57.8 51.2 46.1
+ARL [15] 2019 TAC 51.9 45.9 13.7 79.2 75.5 82.8 0.1 44.9 59.2 47.5 50.1
 […]
-SupHCL [62] 2022 ACM MM 52.5 58.8 70.0 53.5 51.4 73.1 95.6 58.0 64.1
-AAR [60] 2023 TIP 62.4 53.6 71.5 39.0 48.8 76.1 91.3 70.6 64.2
-KDRL [63] 2022 CVPR 60.4 59.2 67.5 52.7 51.5 76.1 91.3 57.7 64.5
-MMA-Net [41] 2023 PRL 63.8 54.8 73.6 39.2 61.5 73.1 92.3 70.5 66.0
-AUNet [64] 2023 Proc.IEEE 60.3 59.1 69.8 48.4 53.0 79.7 93.5 64.7 66.1
-AUNCE(Ours) - 61.8 58.9 74.9 49.7 56.2 73.5 92.1 64.2 66.4
-V-A Comparison with State-of-the-Art Methods
-While our discriminative contrastive learning primarily operates as a self-supervised framework, our experimental results on BP4D and DISFA still compare with some supervised frameworks, because we use AU labels assisting the process of selecting positive sample pairs.
-The focus of our comparison lies in evaluating state-of-the-art methods from the past three years, outlined as follows:
-  • HMP-PS (CVPR 2021) [59] presents a pioneering approach to message passing neural networks incorporating performance-driven structures.
-  • SEV-Net (CVPR 2021) [61] proposes a strategy for AU detection by leveraging semantic embedding and visual features.
+AAR [79] 2023 TIP 66.3 53.9 23.7 81.5 73.6 84.2 43.8 53.8 58.2 46.5 58.5
+MAL [75] 2023 TAC 52.4 57.0 54.1 74.5 78.0 84.9 43.1 47.7 54.4 51.9 59.8
+MoCo [32] 2020 CVPR 35.9 45.4 13.5 83.4 71.3 78.1 23.3 37.3 26.6 50.7 46.5
+SimCLR [22] 2020 ICML 39.6 48.3 5.6 80.7 76.2 80.6 18.1 41.6 46.1 43.8 48.1
+CLP [28] 2023 TIP 44.6 58.7 34.7 75.9 78.6 86.6 20.3 44.8 56.4 42.2 54.3
+AUNCE(Ours) - 53.6 62.8 51.3 80.6 82.1 88.2 48.2 49.8 54.6 58.9 63.0
+(±$\pm 0.33$) (±$\pm 0.18$) (±$\pm 0.28$) (±$\pm 0.24$) (±$\pm 0.41$) (±$\pm 0.11$) (±$\pm 0.35$) (±$\pm 0.23$) (±$\pm 0.42$) (±$\pm 0.35$) (±$\pm 0.29$)
+TABLE V: Quantitative comparison results on Aff-Wild2 dataset, where the evaluation metric is F1-score(%).
+Method Source AU Index Avg.
+1 2 4 6 7 10 12 15 23 24 25 26
+EAC-Net [7] 2018 TPAMI 49.6 33.7 55.6 66.4 82.3 81.4 76.9 11.8 12.5 12.2 93.7 26.8 50.2
+ARL [15] 2019 TAC 59.2 48.2 54.9 70.0 83.4 80.3 72.0 0.1 0.1 17.3 93.0 37.5 51.3
 […]
-  • SupHCL (ACM MM 2022) [62] introduces a method for AU detection that utilizes supervised hierarchical contrastive learning to ensure knowledge consistency across different facial images and AUs.
-  • KDRL (CVPR 2022) [63] employs AU labeling rules from the Facial Action Coding System for self-supervised representation learning in AU detection.
-  • MEGraph (IJCAI 2022) [51] focuses on deep learning of a unique graph to explicitly describe relationships between AUs in the target facial display.
-  • MMA-Net (PRL 2023) [41] proposes a novel AU detection framework based on multiview mixed attention.
-  • AUNet (Proc. IEEE 2023) [64] presents AU detection coupled with facial landmark detection, a simple yet robust baseline for landmark-based AU detection.
-  • MAL (TAC 2023) [58] employs a meta learning approach to automatically select highly related facial expression samples using adaptative weights for training facial expression samples in AU detection.
-  • AAR (TIP 2023) [60] introduces an adaptive attention regression network to regress the global attention map of each AU, guided by attention predefinition and AU detection.
-  • CLP (TIP 2023) [23] aims to alleviate data scarcity issues by learning AU representations from a large pool of unlabeled facial videos in a contrastive learning paradigm.
-The quantitative comparative results for the BP4D dataset are outlined in Table I. AUNCE exhibits superior performance in overall AU detection when contrasted with alternative approaches.
-In comparison to MMA-Net, HMP-PS, AAR, SEV-Net, and KDRL, our approach exhibits a performance advantage of approximately 1.2%-2.3% in F1-score.
-This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing pertinent information for the training process.
-When contrasted with other contrastive learning methods such as CLP and SupHCL, our method demonstrates an average improvement of 3.3% and 1.3%, respectively.
-This improvement is attributed to our positive sample sampling strategy, which alleviates issues associated with AU labels tainted by noise and errors, enabling the encoder to extract differences among static image features more comprehensively.
+PASN [72] 2021 ICCV 65.7 64.2 66.5 76.6 74.7 72.7 78.6 18.5 10.6 55.1 80.7 41.7 58.8
+AAR [79] 2023 TIP 65.4 57.9 59.9 73.2 84.6 83.2 79.9 21.8 27.4 19.9 94.5 41.7 59.1
+AUNCE(Ours) - 68.8 67.6 64.7 74.8 88.6 78.4 81.3 26.9 36.7 18.6 96.1 43.8 62.2
+(±$\pm 0.25$) (±$\pm 0.37$) (±$\pm 0.28$) (±$\pm 0.33$) (±$\pm 0.12$) (±$\pm 0.47$) (±$\pm 0.31$) (±$\pm 0.17$) (±$\pm 0.08$) (±$\pm 0.42$) (±$\pm 0.39$) (±$\pm 0.32$) (±$\pm 0.29$)
+IV-B Comparison with State-of-the-Art Methods
+IV-B1 Evaluation on BP4D
+Since we use AU labels to assist the process of selecting positive sample pairs, we introduce five groups of supervised AU detection frameworks to compare with our method.
+The quantitative results of state-of-the-art methods from the past seven years for BP4D dataset are shown in Table II.
+AUNCE exhibits superior performance in AU detection when contrasted with alternative approaches.
+Methods in the first group are based on feature extraction of regional area around AUs, including EAC-Net [7], ROI [4] and ARL [15].
+Our approach achieves a notable improvement, with an increase in F1-score ranging from 5.0% to 10.2% over these methods.
+This advancement is primarily due to early feature extraction methods neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU.
+Furthermore, the precision of the encoders utilized by these methods is generally lower.
 […]
-Compared to the second-best approach, MEGraph, our method demonstrates superior performance by elevating the F1-score from 65.5% to 65.7%.
-It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly half.
+Methods in the third group mainly utilize the AU semantic prior knowledge, including MMA-Net [76] and SEV-Net [77].
+Our approach demonstrates an average improvement of 2.7% and 2.2%, respectively.
+This phenomenon can be attributed to the inherent limitations of these methods in effectively capturing and representing the distinct information encoded within Action Units (AUs) using simple semantic descriptions.
+Methods in the fourth group focus on capturing correlational information among AUs, including HMP-PS [78], AAR [79], ISTR [80] and MEGraph [56].
+Our approach exhibits a performance advantage of approximately 0.6%-2.7% in F1-score.
+This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing differential information instead of pixel-level information.
+It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly 7M.
 This highlights the applicability of AUNCE to more lightweight network structures and emphasizes that learning only the difference information among features is more potent than learning the pixel-level information of the entire image.
-Comparable outcomes are evident in the DISFA dataset, as outlined in Table II, where AUNCE attains the highest overall performance in AU detection.
-V-B Ablation Study
-TABLE III: Quantitative results of ablation study on BP4D dataset.
-Model BL PS IR F1-Score(%) Accuracy(%)
-A $\checkmark$ 60.1 79.1
-B $\checkmark$ $\checkmark$ 63.2(+3.1) 79.7(+0.6)
-C $\checkmark$ $\checkmark$ 63.4(+3.3) 79.8(+0.7)
-D $\checkmark$ $\checkmark$ $\checkmark$ 65.7(+5.6) 80.2(+1.1)
-V-B1 Ablation Study
-The quantitative results of our method with various component combinations are summarized in Table III, encompassing the baseline (BL) rooted in the new discriminative contrastive learning paradigm, utilizing images with the highest similarity to original images as positive sample pairs, the positive sample sampling strategy (PS), and the importance re-weighting strategy to minority AUs (IR).
-The merits of our proposed discriminative contrastive learning paradigm will be elaborated in Section V-B2.
-Contribution of the positive sample sampling strategy: As outlined in Table III, the positive sample sampling strategy contributes to a 3.1% increase in F1-score and a 0.6% improvement in Accuracy compared to the baseline.
+Methods in the last group are based on contrastive learning, including SIMCLR [22], MoCo [32], CLP [28], BTFM [81], SupHCL [49], KSRL [46] and CLEF [82].
+When compared to these methods, our approach demonstrates a performance advantage of approximately 0.2% to 16.3% in F1-score.
+This improvement is attributed to the efficacy of our proposed negative sample re-weighting strategy and positive sample sampling strategy, specifically tailored for the AU detection task using static images.
+IV-B2 Evaluation on DISFA
+Comparable outcomes are evident in the DISFA dataset, as outlined in Table III, where AUNCE attains the highest overall performance in AU detection.
+Furthermore, a comparison between Table II and Table III reveals that recent methods such as SEV-Net, MEGraph, and BTFM demonstrate strong performance on the BP4D dataset but achieve only mediocre results on the DISFA dataset.
+This disparity suggests their limited generalizability.
+The more pronounced class imbalance issue in DISFA, as shown in Table I, likely contributes to this performance gap.
+In contrast, AUNCE consistently delivers robust performance on both BP4D and DISFA datasets and achieves more stable results across various AUs due to our proposed negative sample re-weighting strategy.
+IV-B3 Evaluation on GFT
+Table IV presents the F1-score results on the GFT dataset, where our AUNCE method demonstrates significantly superior performance compared to previous approaches.
+Unlike BP4D and DISFA, which primarily consist of near-frontal face images, GFT includes many images captured under out-of-plane poses.
+In this challenging scenario, other methods such as EAC-Net, ARL and CLP so on struggle to perform well.
+In contrast, AUNCE achieves a commendable average F1-score of 63.0%.
+IV-B4 Evaluation on Aff-Wild2
+[…]
+IV-C Ablation Study
+IV-C1 Compared with weighted cross-entropy loss
+TABLE VI: Comparison of the weighted cross-entropy (WCE) loss and AUNCE with different encoders on BP4D dataset.
+Backbone Source Loss F1-Score(%) Accuracy(%)
+ResNet18 [88] 2016CVPR WCE 52.3 73.8
+AUNCE $\textbf{54.9}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.6\uparrow}}$ $\textbf{74.3}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.5\uparrow}}$
+ResNet50 [88] 2016CVPR WCE 56.2 75.5
+AUNCE $\textbf{59.7}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3.5\uparrow}}$ $\textbf{77.0}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.5\uparrow}}$
+ROI [4] 2017CVPR WCE 56.4 75.9
+AUNCE $\textbf{58.8}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.4\uparrow}}$ $\textbf{76.7}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.8\uparrow}}$
+ARL [15] 2019TAC WCE 61.1 78.2
+AUNCE $\textbf{62.9}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.8\uparrow}}$ $\textbf{79.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.9\uparrow}}$
+[…]
+AUNCE $\textbf{64.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.7\uparrow}}$ $\textbf{80.2}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.6\uparrow}}$
+MEGraph [56] 2022IJCAI WCE 63.6 80.0
+(Onestage) AUNCE $\textbf{66.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.5\uparrow}}$ $\textbf{80.4}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.4\uparrow}}$
+Due to the primary innovation of this paper being the introduction of a novel contrastive loss, AUNCE, which is compatible with various AU detection encoders, we conducted a comparative analysis between the proposed AUNCE loss and the traditional weighted cross-entropy loss (WCE) across multiple AU detection encoders, as detailed in Table VI.
+The findings reveal that AUNCE demonstrates superior performance in AU detection.
+Notably, the MEGraph encoder achieved the best results, with the F1-score on the BP4D dataset improving from 63.6% to 66.1%.
+This underscores the effectiveness of capturing unique features of different samples over focusing solely on pixel-level information.
+Furthermore, we optimized the two-stage training process of MEGraph into a single-stage process, significantly reducing both the training time and the number of parameters required for the model.
+TABLE VII: Quantitative results of ablation study on BP4D dataset.
+Model BL wi PS NS AU Index F1-Score(%) Accuracy(%)
+1 2 4 6 7 10 12 14 15 17 23 24
+A $\checkmark$ 39.8 44.3 58.8 76.0 77.7 76.9 88.0 64.5 49.0 59.0 43.5 43.4 60.1 79.1
+B $\checkmark$ $\checkmark$ 40.1 43.0 58.9 75.4 79.3 78.8 87.8 64.1 48.7 61.0 46.9 47.5 61.00.9↑ 79.30.2↑
+C $\checkmark$ $\checkmark$ $\checkmark$ 51.1 45.6 58.3 77.5 77.0 84.5 89.0 63.7 54.6 65.5 46.8 53.3 63.83.7↑ 79.80.7↑
+D $\checkmark$ $\checkmark$ $\checkmark$ 50.2 43.0 58.7 77.9 77.3 83.7 88.7 64.2 53.9 64.9 50.0 56.4 64.14.0↑ 79.90.8↑
+E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 53.6 49.8 61.6 78.4 78.8 84.7 89.6 67.4 55.1 65.4 50.9 58.0 $\textbf{66.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}6.0\uparrow}}$ $\textbf{80.4}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.3\uparrow}}$
+IV-C2 Evaluating other components
+We further evaluate the effectiveness of weight coeffcient $\omega_{i}$, negative sample re-weighting strategy and positive sample samplying strategy.
+Quantitative results of our method with the above component combinations are summarized in Table VII.
+The baseline (BL) is rooted in SIMCLR [22], but positive sample pairs are changed to images with the highest similarity to original images.
+Other components are the weight coefficient for both WCE and AUNCE ($w_{i}$), the positive sample sampling strategy (PS), and the negative sample re-weighting strategy (NS).
+Weight coefficient $w_{i}$: The weights $w_{i}$ have been pointed out as a fixed value in Section.
+III-A, which achieve an increase of 0.9% in F1-score and 0.2% in Accuracy.
+For an AU, its activation indicator is a binary variable, and the corresponding information content can be measured using entropy: $-\log(p_{i})$.
+Therefore, AUs with lower activation probabilities carry more information and should be assigned higher weights.
+This forms the theoretical basis for our weighting strategy.
+Positive sample sampling strategy: As outlined in Table VII, the positive sample sampling strategy contributes to a 2.8% increase in F1-score and a 0.5% improvement in Accuracy compared to model B.
 This validates that selecting different types of positive sample pairs with appropriate probabilities can mitigate issues arising from AU labels tainted by noise and errors.
@@
 These pairs provide the models with diverse information from various views, thereby enhancing the generalization ability of the contrastive learning framework.
-Effectiveness of the importance re-weighting strategy: The proposed importance re-weighting strategy to minority AUs results in a 3.3% increase in F1-score compared to model A.
-This illustrates that the strategy can change the backpropagation gradient of positive and negative samples, facilitating faster label category gradient updates with a smaller number of samples.
-Additionally, model D achieves an average increase of 2.5% in F1-score compared to model B, further supporting the efficacy of the importance re-weighting strategy.
-Based on ablation studies, the introduction of each component proves beneficial for performance improvement, with IR emerging as the most effective component.
-TABLE IV: Comparison of the weighted cross-entropy loss and AUNCE on BP4D and DISFA.
-Dataset Loss MEGraph-stage1
-F1-Score Accuracy
-BP4D WCE 63.6 80.0
-AUNCE 65.7(+2.1) 80.2(+0.2)
-DISFA WCE 62.2 93.7
-AUNCE 66.4(+4.2) 94.3(+0.6)
-V-B2 Compared with weighted cross-entropy loss
-As depicted in Table IV, AUNCE has exhibited superior performance compared to the traditional weighted cross-entropy loss (WCE) for AU detection, resulting in an increase in F1-score from 63.6% to 65.7% on BP4D and from 62.2% to 66.4% on DISFA.
-This highlights that capturing the distinctive features of different samples proves more effective than concentrating on pixel-level information.
-Consequently, our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance problem for each AU type, and bolsters model robustness against AU labels affected by noise and errors.
+Negative sample re-weighting strategy: The negative sample re-weighting strategy to minority AUs results in a 3.1% increase in F1-score and a 0.6% increase in Accuracy compared to model B.
+This illustrates that the strategy can change the backpropagation gradient of minority and majority class samples, facilitating faster label category gradient updates with a smaller number of samples.
+Additionally, model E achieves an average increase of 2.3% in F1-score compared to model C, further supporting the efficacy of the re-weighting strategy.
+Figure 4: T-SNE visualization of feature representations on BP4D and DISFA datasets.
+Colors indicates whether AU1 exists.
+Top row: (a) Representations optimized by WCE on BP4D dataset, and (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E.
+Bottom row: (g) Representations optimized by AUNCE on DISFA dataset, and (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.
+We further show the visualization using T-SNE dimensionality reduction of the learned feature representations optimized by AUNCE and WCE to qualitatively assess the efficacy of our proposed discriminative contrastive learning paradigm, as illustrated in Fig.
+4.
+For each dataset, a random sample of 200 images is utilized for illustration purposes.
+A comparison between Fig.
+4(a), Fig.
+4(f) against Fig.
+4(g), Fig.
+4(l) reveals that the learned representations optimized by AUNCE exhibit greater separability than those optimized by WCE.
+Additionally, Fig.
+4(b)-(f) and Fig.
+4(g)-(l) qualitatively showcase the effectiveness of each component in our ablation study.
+Based on all ablation studies, it demonstrates that our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance issue of each AU type, and bolsters model robustness against AU labels affected by noise and errors.
 This, in turn, leads to enhanced generalization performance in AU detection.
-AUNCE can seamlessly integrate into various lightweight AU detection networks, serving as a replacement for generative models.
-The loss demonstrates robust compatibility in diverse model architectures.
-V-C Hyperparameters Analysis
-TABLE V: Quantitative results the probabilities of various positive sample pairs $p_{1},p_{2},p_{3},p_{4}$ on BP4D dataset.
-$p_{1}$ $p_{2}$ $p_{3}$ $p_{4}$ F1-Score
+The introduction of each component proves beneficial for performance improvement, with NS emerging as the most effective component.
+Moreover, the results on each AU are also shown in Table VII.
+It can be seen that the results on basically every AU have improved or approached after we add $w_{i}$, PS and NS.
+IV-D Hyperparameters Analysis
+TABLE VIII: Quantitative results the probabilities of various positive sample pairs $p_{1},p_{2},p_{3},p_{4}$ on BP4D dataset.
+$p_{1}$ $p_{2}$ $p_{3}$ $p_{4}$ F1-Score(%)
 Probabilities 1 0 0 0 60.1
-0 1 0 0 42.0
+0 1 0 0 48.4
 0 0 1 0 63.2
@@
 0.1 0.1 0.8 0 63.8
+0.15 0.15 0.7 0 66.1
 0.2 0.2 0.6 0 65.7
+0.25 0.25 0.5 0 65.2
 0.3 0.3 0.4 0 64.3
 0.4 0.4 0.2 0 58.8
-V-C1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
+IV-D1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
 In our paper, $p_{1}$ represents the probabilities of images with the highest similarity to original images, $p_{2}$ represents the probabilities of self-supervised images enhanced by simple augmentations, $p_{3}$ represents the probabilities of a mixture of all positive samples, and $p_{4}$ represents the probabilities of images with the lowest similarity to original images.
@@
 As $p_{1}$ and $p_{2}$ are deemed reliable positive samples, we set their values to be the same during the variation.
-The experimental results are presented in Table V.
-When $p_{1}$=$p_{2}$=0.2, $p_{3}$=0.6, and $p_{4}$=0, the experimental result is optimal.
-It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent errors and noise.
+The experimental results are presented in Table VIII.
+When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal.
+It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors.
+The conclusion is consistent with  [89, 90].
 The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class.
-The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of error and noisy samples, thereby reinforcing the robustness of the training process.
-Refer to caption (a)
-Refer to caption (b)
+The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.
 Figure 5: Quantitative results of $\beta$ of different positive and negative samples for each AU type on BP4D dataset.
-From up to down: (a) $\beta$ for positive samples (the minority class), and (b) $\beta$ for negative samples (the majority class).
-V-C2 Backpropagation rate controller $\beta$
-As is well understood, the hyperparameter $\beta$ plays a crucial role in governing the relative importance of positive and negative samples during the parameter update process.
+From left to right: (a) $\beta$ for positive samples (the minority class), and (b) $\beta$ for negative samples (the majority class).
+IV-D2 Backpropagation rate controller $\beta$
+As is well understood, the hyperparameter $\beta$ plays a crucial role in governing the relative importance of minority and majority class samples during the parameter update process.
 It ensures that the model allocates more attention to minority class samples when necessary and balances the gradient magnitudes accordingly.
@@
 The experimental results are depicted in Fig.
-3.
+5.
 The optimal experimental outcome is observed when $\beta$ for the minority class is set to 1.2, and $\beta$ for the majority class is fixed at 0.4.
-Hence, when $\beta_{\text{minority}}>\beta_{\text{majority}}$, the impact of the imbalanced problem can be effectively mitigated.
-VI Conclusion
-In this paper, we introduce an innovative discriminative contrastive learning paradigm designed to replace traditional pixel-level learning in the domain of AU detection.
-Our proposed contrastive loss, referred to as AUNCE, shares similarities with InfoNCE, transforming the task of learning image data distribution information into the more effective learning of differentiated information among images.
+Hence, when $\beta_{\text{minority}}>\beta_{\text{majority}}$, the impact of the class imbalance issue can be effectively mitigated.
+V Conclusion
+In this paper, we introduce a contrastive learning framework for AU detection.
+Our proposed contrastive loss, named AUNCE, shares similarities with InfoNCE, transforming the task of learning entire pixel-level information of AUs into the more effective learning of subtle feature differences when AUs activate.
 Additionally, we present a positive sample sampling strategy that introduces three types of positive sample pairs, addressing the challenge of AU labels tainted by noise and errors in two AU datasets.
-To further enhance the framework, we incorporate a importance re-weighting strategy specifically tailored to minority AUs.
-This strategy alters the backpropagation gradient of positive and negative samples, effectively addressing the imbalanced nature of each AU type.
+To further enhance the framework, we incorporate a negative sample re-weighting strategy specifically aimed for minority AUs.
+This strategy controls the relative contribution of minority class samples to the parameter updates, effectively addressing the class imbalance issue of each AU type, and contributing to hard negative sample mining.
 Experimental results underscore the superior performance of AUNCE in comparison to existing methods.
 Acknowledgement
+This research was supported by the National Natural Science Foundation of China (Nos.
+62176221, 61572407), and Sichuan Science and Technology Program (Nos.
+2024NSFTD0036, 2024ZHCG0166).
 The computation is completed in the HPC Platform of Huazhong University of Science and Technology.